### v1

# Controllability score for linear time-invariant systems on an infinite time horizon

Abstract
We introduce a scaled controllability Gramian that can be computed reliably even for unstable systems. Using this scaled Gramian, we reformulate the controllability scoring problems into equivalent but numerically stable optimization problems. Their optimal solutions define dynamics-aware network centrality measures, referred to as the volumetric controllability score (VCS) and the average energy controllability score (AECS). We then formulate controllability scoring problems on an infinite time horizon. Under suitable assumptions, we prove that the resulting VCS and AECS are unique and that the finite-horizon scores converge to them. We further show that VCS and AECS can differ markedly in this limit, because VCS enforces controllability of the full system, whereas AECS accounts only for the stable modes. Finally, using Laplacian dynamics as a representative example, we present numerical experiments that illustrate this convergence.
Index Terms:
Network centrality, controllability score, infinite horizon, linear time-invariant system
I Introduction
Large-scale networks play an essential role in understanding complex phenomena and in intervening in complex dynamical systems across diverse domains [19]. Many systems of contemporary interest—such as brain networks [4, 29], biological interaction networks [2], social networks [1], and power grids [5, 20]—are represented as networks consisting of a large number of interacting components. In such systems, macroscopic behavior frequently emerges from collective interactions rather than from individual units alone. While structural descriptions are informative, they are frequently insufficient to predict or steer real-world network dynamics, motivating frameworks that explicitly incorporate dynamics in addition to topology [16, 17, 26].
Control theory, and in particular controllability, provides a dynamics-aware framework for analyzing and intervening in complex networked systems [16, 17]. Controllability characterizes whether a system can be steered to desired states [13] and leads to quantitative descriptors that reflect both dynamics and interconnection structure. Controllability-based analyses have proved useful in neuroscience [9] and biology [28], among others. These developments motivate node-level measures that quantify where to intervene and how strongly, while accounting for the control effort required to influence the network.
In this context, controllability scores [23] provide controllability-based centrality measures for continuous-time linear time-invariant (LTI) dynamical networks of the form
$\dot{x}(t)=Ax(t),$ (1)
where $x(t)\in\mathbb{R}^{n}$ is the state vector collecting the states of $n$ nodes, and the system matrix $A=(a_{ij})\in\mathbb{R}^{n\times n}$ captures weighted interactions: $a_{ij}$ represents the influence of node $j$ on node $i$. We use “node ii” and “state $x_{i}$” interchangeably. Under mild conditions, controllability scores are defined as the unique optimal solutions of convex optimization problems, thereby assigning each node a quantitative measure of its importance for controlling the network dynamics.
The controllability scores are designed to address limitations of existing controllability-based frameworks. Prior approaches can be broadly categorized into structural controllability [15] and quantitative controllability measures [18]. Structural controllability depends only on topology and enables efficient graph-theoretic analysis without identifying edge weights [16, 17], but it does not preclude prohibitively large control-energy requirements [30]. Quantitative measures incorporate edge weights and time horizons to capture energy-related aspects of controllability; building on this viewpoint, controllability scores evaluate node importance while accounting for control effort. Existing Gramian-based centralities such as volumetric controllability centrality (VCE) and average control energy centrality (ACE) may still be misaligned with quantitative controllability because they rely on the rank or pseudoinverse of a single-input controllability Gramian [23]. By contrast, controllability scores are defined by introducing virtual actuation at all state nodes and optimizing the input allocation, rather than relying on a single-input controllability Gramian. Subsequent studies established uniqueness properties [23, 22], demonstrated applicability to human brain networks [22], and incorporated constraints on actuated nodes [24].
In Section II-A, we define two controllability scores (VCS and AECS) as optimal solutions to distinct controllability-Gramian–based optimization problems. This finite-horizon formulation requires specifying a terminal time $T>0$. However, the choice of $T$ has not been systematically addressed in the existing literature. It has been reported in [22] that when $T$ is small, the resulting scores are dominated by short-term dynamics and fail to capture meaningful system-wide properties, whereas for sufficiently large $T$, they reflect the network-structured dynamics of the system. These observations suggest taking $T$ large. A natural extreme is $T=\infty$, which eliminates dependence on the system time scale and on the researcher’s subjective choice, and is therefore desirable from the viewpoint of reproducibility. However, except for the case where $A$ is Hurwitz (i.e., all its eigenvalues have strictly negative real parts), the controllability Gramian diverges as $T\to\infty$, so the scores are no longer well defined. Moreover, even for large but finite $T$, the Gramian can become ill-conditioned, making objective-function evaluation numerically unstable and potentially preventing accurate computation. This paper provides a solution to this issue.
The contributions of this paper are summarized as follows.
  1. 1.
    We introduce a scaling method for the controllability Gramian that compensates for the contributions of unstable modes of $A$ and ensures convergence as $T\to\infty$. We then reformulate the controllability scores using the scaled Gramian, which yields equivalent optimization problems but enables numerically stable computation for large values of $T$.
  2. 2.
    By taking the limit $T\to\infty$ in the reformulated problems, we obtain controllability scores on an infinite time horizon, and prove, under suitable assumptions, that the finite-horizon scores converge to them. This provides a principled choice of the terminal time, namely $T=\infty$, and also offers an explanation for the empirical observation reported in [22] that, although the controllability Gramian diverges for unstable systems, the controllability scores nevertheless converge to a certain value as $T\to\infty$. This infinite-horizon formulation also highlights an important advantage of defining centrality through optimization problems. While several centrality measures based on functions of the controllability Gramian have been proposed [25], including trace- or log-determinant–based ones, these quantities are generally not well-defined on an infinite time horizon unless the system matrix $A$ is Hurwitz. This is because the controllability Gramian diverges as $T\to\infty$, as discussed in Remark 1. In contrast, since the controllability scores are defined as the solutions to optimization problems over a bounded feasible set, the scores themselves can remain meaningful even when the controllability Gramian diverges.
  3. 3.
    We theoretically show that VCS and AECS remain markedly different on the infinite time horizon, as in the finite-horizon case reported in [22]. In particular, the infinite-horizon AECS can yield an allocation at which the associated controllability Gramian becomes singular, reflecting that AECS accounts only for controllability of the stable modes, whereas VCS enforces controllability of the entire system. We illustrate these phenomena using Laplacian dynamics as a representative example.
  4. 4.
    Under suitable assumptions, we show that the scaled controllability Gramian can be computed in a numerically stable manner using a real Schur decomposition and the solution of a Sylvester equation. This enables the stable numerical evaluation of the infinite-horizon controllability scores.
The rest of this paper is organized as follows. In Section II, we summarize the controllability scores and clarify remaining challenges. In Section III, we describe a scaling method for the controllability Gramian that ensures convergence as $T\to\infty$, and reformulate the controllability scores using the scaled controllability Gramian. In Section IV, we formulate the infinite-horizon controllability scores and prove, under suitable assumptions, that they are uniquely determined and that the finite-horizon controllability scores converge to them. In Section V, we present a numerically stable algorithm for computing the infinite-horizon scaled controllability Gramian and review an existing optimization algorithm for computing the infinite-horizon controllability scores. In Section VI, we present numerical experiments demonstrating the convergence of the finite-horizon controllability scores to their infinite-horizon counterparts. Finally, Section VII concludes this paper.
Notation
Throughout this paper, $n$ denotes the number of nodes in the network, and $A\in\mathbb{R}^{n\times n}$ denotes the system matrix in 1. The sets of all natural numbers, all real numbers, and all complex numbers are denoted by $\mathbb{N}$, $\mathbb{R}$, and $\mathbb{C}$, respectively. Let $e_{i}\coloneq(0,\ldots,0,1,0,\ldots,0)^{\top}\in\mathbb{R}^{n}$ be a standard vector whose $i$-th entry is $1$ and other entries are 0. Let $I_{k}$ and $N_{k}$ denote the identity matrix of size $k$ and the nilpotent Jordan block of size $k$ with ones on the superdiagonal, respectively. Let $O$ denote a zero matrix of a suitable size.
Let $\Delta$ and $\widetilde{\Delta}$ be the standard simplices in $\mathbb{R}^{n}$ and $\mathbb{R}^{n-1}$, respectively, i.e.,
$\Delta\coloneq\left\{(p_{i})\in\mathbb{R}^{n}\,\middle|\,\begin{gathered}p_{i}\geq 0\ (i=1,\ldots,n),\\ \textstyle\sum_{i=1}^{n}p_{i}=1\end{gathered}\right\},$
$\widetilde{\Delta}\coloneq\left\{(\widetilde{p}_{i})\in\mathbb{R}^{n-1}\,\middle|\,\begin{gathered}\widetilde{p}_{i}\geq 0\ (i=1,\ldots,n-1),\\ \textstyle\sum_{i=1}^{n-1}\widetilde{p}_{i}\leq 1\end{gathered}\right\}.$
The mapping $\varphi\colon\widetilde{\Delta}\to\Delta$ defined by
$\varphi(\widetilde{p}_{1},\ldots,\widetilde{p}_{n-1})=(\widetilde{p}_{1},\ldots,\widetilde{p}_{n-1},1-\sum_{i=1}^{n-1}\widetilde{p}_{i})^{\top}$
is a homeomorphism; its inverse is given by
$\varphi^{-1}(p_{1},\ldots,p_{n})=(p_{1},\ldots,p_{n-1})^{\top}.$
For $\lambda\in\mathbb{C}$, $\operatorname{Re}(\lambda)$, $\operatorname{Im}(\lambda)$, and $\overline{\lambda}$ denote the real part, the imaginary part, and the complex conjugate of $\lambda$, respectively. For a real matrix $\in\mathbb{R}^{k\times\ell}, XX^{\top}$ denotes the transpose of $X$, and for a complex matrix $\in\mathbb{C}^{k\times\ell}, YY^{*}$ denotes the conjugate transpose of $Y$. For a real invertible matrix $X\in\mathbb{R}^{\ell\times\ell}$, we write $X^{-\top}\coloneq(X^{-1})^{\top}$, and for a complex invertible matrix $Y\in\mathbb{R}^{\ell\times\ell}$, we write $Y^{-*}\coloneq(Y^{-1})^{*}$. For a matrix $Z\in\mathbb{C}^{\ell\times\ell}$, $\mathrm{e}^{Z},\det Z$, and $\operatorname{tr}Z$ denote the exponential of $Z$, the determinant of $Z$, and the trace of $Z$. For $C=(c_{ij})\in\mathbb{R}^{\ell\times\ell}$ and $^{\prime}\in\mathbb{R}^{\ell^{\prime}\times\ell^{\prime}}, CC$$\otimes C^{\prime}\coloneq\begin{pmatrix}c_{11}C^{\prime}&\cdots&c_{1\ell}C^{\prime}\\ \vdots&\ddots&\vdots\\ c_{\ell 1}C^{\prime}&\cdots&c_{\ell\ell}C^{\prime}\end{pmatrix}\in\mathbb{R}^{\ell\ell^{\prime}\times\ell\ell^{\prime}}$ denotes their Kronecker product. For Hermitian matrices $X,Y\in\mathbb{C}^{\ell\times\ell}$, we write $X\succ Y$ if XYX-Y is positive definite. For $v_{1},\ldots,v_{\ell}\in\mathbb{C}$, let $\mathrm{diag}(v_{1},\ldots,v_{\ell})$ denote the diagonal matrix with the diagonal elements $v_{1},\ldots,v_{\ell}$. For a vector $v=(v_{1},\ldots,v_{\ell})^{\top}\in\mathbb{C}^{\ell}$, $\left\|{v}\right\|\coloneq\sqrt{\sum_{i=1}^{\ell}|v_{i}|^{2}}$ denotes the standard Euclidean norm. For a matrix $X\in\mathbb{C}^{k\times\ell}$, $\left\|{X}\right\|\coloneq\sqrt{\operatorname{tr}(X^{*}X)}$ denotes the Frobenius norm.
II Preliminaries
II-A Controllability score
In this subsection, we summarize the concept of the controllability scores [23]. To define them, we introduce a virtual input allocation in which a fixed amount of control resources is distributed across network nodes. Let $p\in\Delta$ and consider the following system illustrated in Fig. 1:
$\dot{x}(t)=Ax(t)+\operatorname{diag}(\sqrt{p_{1}},\ldots,\sqrt{p_{n}})u(t),$ (2)
where $u(t)\coloneq(u_{1}(t),\ldots,u_{n}(t))^{\top}\in\mathbb{R}^{n}$ is a virtual control signal, and each node $x_{i}$ is actuated by $\sqrt{p_{i}}u_{i}(t)$. The value $p_{i}$ represents the amount of control resources allocated to node $x_{i}$; under a constraint on the input energy, a larger value of $p_{i}$ allows a stronger control effect on node $x_{i}$, indicating that node $x_{i}$ is a central node in the system. The controllability score associates each coefficient $p_{i}$ with the importance of node $x_{i}$.
$x_{1}$$x_{2}$$x_{3}$$x_{4}$$x_{5}$$u_{1}$$u_{2}$$u_{3}$$u_{4}$$u_{5}$$\sqrt{p_{1}}$$\sqrt{p_{2}}$$\sqrt{p_{3}}$$\sqrt{p_{4}}$$\sqrt{p_{5}}$
Figure 1: Virtual system for defining controllability scores. The node $x_{i}$ denotes a state variable, and the node $u_{i}$ denotes a virtual control signal. The coefficients $p_{i}$ specify how a fixed amount of control resources is allocated to the nodes, and each input $u_{i}$ corresponds one-to-one with node $x_{i}$ through the weight $\sqrt{p_{i}}$.
Let us maximize the controllability of system 2 with respect to $p$ for a certain measure, and let $p^{*}$ denote the optimal solution. The optimal solution $p^{*}$ represents the allocation of control authority that best enhances system controllability. The entries $p_{i}^{*}$ of this allocation, therefore, quantify how important each node $x_{i}$ is for controlling the system dynamics. Accordingly, $p^{*}$ can be regarded as a centrality measure, and we refer to it as a controllability score.
The controllability of a system can be measured by the controllability Gramian. For $B\in\mathbb{C}^{n\times n}$, let $\widetilde{W}(B;T)\in\mathbb{C}^{n\times n}$ be defined as follows:
$\widetilde{W}(B;T)\coloneq\int_{0}^{T}\mathrm{e}^{At}B\mathrm{e}^{A^{\top}t}\mathrm{d}t.$ (3)
The controllability Gramian for 2 with terminal time $T>0$ is given by $\widetilde{W}(\operatorname{diag}(p);T)$, which we abbreviate as $\widetilde{W}(p;T)$. The quantities $\log\det\widetilde{W}(p;T)$ and $\operatorname{tr}\left(\widetilde{W}(p;T)^{-1}\right)$ can serve as scalar controllability measures. A smaller value of $\log\det\widetilde{W}(p;T)$ corresponds to a larger volume of the reachable set $\mathcal{E}(p;T)$, defined as
[…]
A smaller value of $\operatorname{tr}\left(\widetilde{W}(p;T)^{-1}\right)$ implies a smaller average minimum energy to steer the state from the origin to points on the unit sphere $\{y\in\mathbb{R}^{n}\mid\left\|{y}\right\|=1\}$, where the average is taken with respect to the uniform distribution on the sphere.
We thus consider the following minimization problems, which we refer to as controllability scoring problems.
Problem 1
$\operatorname*{minimize}_{p}\quad$ ${-}\log\det\widetilde{W}(p;T)$
$\mathrm{subject\ to}\quad$ $p\in\Delta,\ \widetilde{W}(p;T)\succ O.$
Problem 2
$\operatorname*{minimize}_{p}\quad$ $\operatorname{tr}\left(\widetilde{W}(p;T)^{-1}\right)$
$\mathrm{subject\ to}\quad$ $p\in\Delta,\ \widetilde{W}(p;T)\succ O.$
The optimal solution to Problem 1 is referred to as the volumetric controllability score (VCS), whereas that of Problem 2 is referred to as the average energy controllability score (AECS). As explained above, these quantities can be interpreted as measures of the importance of individual nodes from the perspective of controllability.
II-B Remaining challenges and our approach
While several topics related to the controllability score have been studied [23, 22, 24], the following two challenges remain unresolved as described in Section I:
  1. i)
    It remains unclear how to choose the terminal time $T$. An appropriate selection of $T$ has not been explicitly provided in the existing literature, except in the case where $A$ is Hurwitz, for which the choice $T=\infty$ is admissible [23].
  2. ii)
    A reliable algorithm for unstable systems is lacking. If the matrix $A$ is not Hurwitz, $\widetilde{W}(p;T)$ diverges as $T\to\infty$, which makes the computation of the score for sufficiently large $T$ numerically unstable.
To resolve i) and ii), we introduce a scaling approach for the controllability Gramian $\widetilde{W}(p;T)$. By scaling the components of $\widetilde{W}(p;T)$ associated with the unstable modes of $A$, we define a new matrix $W(p;T)$ that converges as $T\to\infty$. Using $W(p;T)$, we reformulate Problems 1 and 2 into equivalent but numerically stable problems. This reformulation resolves ii). Furthermore, formally taking the limit $T\to\infty$ leads to controllability scoring problems on an infinite time horizon. While such problems have been studied for Hurwitz matrices [23], they have not been formulated for non-Hurwitz matrices. Under certain assumptions, their optimal solutions can be interpreted as the limit of the scores as $T\to\infty$. Therefore, this yields the principled choice of the terminal time $T=\infty$, thereby providing a solution to i).
Remark 1
Note that centrality measures based on functions of the controllability Gramian are not necessarily well defined in the limit $T\to\infty$. For example, average controllability centrality (AC), VCE, and ACE proposed in [25] were originally defined for stable systems, but can be naively extended to unstable systems, in which case they are defined for node $i$ by
$\mathrm{AC}_{i,T}\coloneq\operatorname{tr}\left(\widetilde{W}(e_{i}e_{i}^{\top};T)\right),$
$\mathrm{VCE}_{i,T}\coloneq\sum_{j=1}^{k_{i}}\log\lambda_{j}\left(\widetilde{W}(e_{i}e_{i}^{\top};T)\right),$
$\mathrm{ACE}_{i,T}\coloneq-\sum_{j=1}^{k_{i}}\dfrac{1}{\lambda_{j}\left(\widetilde{W}(e_{i}e_{i}^{\top};T)\right)},$
where $k_{i}\coloneq\operatorname{rank}\left(\widetilde{W}(e_{i}e_{i}^{\top};T)\right)$ and $\lambda_{j}\left(\widetilde{W}(e_{i}e_{i}^{\top};T)\right)\ (j=1,\ldots,k_{i})$ denote the positive eigenvalues. Since eigenvalues of $\widetilde{W}(e_{i}e_{i}^{\top};T)$ that are close to zero reduce the values of $\mathrm{VCE}_{i,T}$ and $\mathrm{ACE}_{i,T}$, whereas zero eigenvalues do not, it has been pointed out in [23] that VCE and ACE do not necessarily provide appropriate centrality measures from the perspective of controllability.
Furthermore, as discussed in [22], capturing the influence of network structure requires choosing the terminal time $T$ to be sufficiently large, which naturally leads one to consider the limit $T\to\infty$. Since $\left\|{\widetilde{W}(e_{i}e_{i}^{\top};T)}\right\|\to\infty$ as $T\to\infty$ for at least one $i$, the maximum eigenvalue of $\widetilde{W}(e_{i}e_{i}^{\top};T)$ also diverges, implying $\mathrm{AC}_{i,T}\to\infty$ and $\mathrm{VCE}_{i,T}\to\infty$.
In contrast, VCS and AECS are defined as optimal solutions to Problems 1 and 2, which is defined over a bounded feasible set. Consequently, even if the controllability Gramian diverges as $T\to\infty$, VCS and AECS themselves can converge to a meaningful value, as discussed in Section IV-B.
III Scaling method of the controllability Gramian and reformulation of Problems 1 and 2
In this section, we reformulate Problem 1 as
Problem 3
$\operatorname*{minimize}_{p}\quad$ ${-}\log\det W(p;T)$
$\mathrm{subject\ to}\quad$ $p\in\Delta,\ W(p;T)\succ O,$
and Problem 2 as
Problem 4
$\operatorname*{minimize}_{p}\quad$ $\operatorname{tr}\left(Q^{-*}D(T)^{-*}W(p;T)^{-1}D(T)^{-1}Q^{-1}\right)$
$\mathrm{subject\ to}\quad$ $p\in\Delta,\ W(p;T)\succ O.$
Here, $W(p;T),Q$, and $D(T)$ are defined in Section III-A. We also prove that $W(p;T)$ converges as $T\to\infty$, and discuss the advantages of Problems 3 and 4.
III-A Problem settings
In this section, the matrix $A\in\mathbb{R}^{n\times n}$ is arbitrary, and we assume that it admits a Jordan decomposition
$Q^{-1}AQ=J\coloneq\begin{pmatrix}J_{1}\\ &\ddots\\ &&J_{m}\end{pmatrix}$
for some invertible matrix $Q\in\mathbb{C}^{n\times n}$. Here, $m$ is the number of Jordan blocks, and $J_{i}=\lambda_{i}I_{n_{i}}+N_{n_{i}}\in\mathbb{C}^{n_{i}\times n_{i}}$ denotes a Jordan block of size $n_{i}$ associated with the eigenvalue $\lambda_{i}$ for $i=1,\ldots,m$. Let $m_{-},m_{0}$, and $m_{+}$ denote the number of Jordan blocks satisfying $\operatorname{Re}(\lambda_{i})<0,\operatorname{Re}(\lambda_{i})=0$, and $\operatorname{Re}(\lambda_{i})>0$, respectively. Moreover, let the eigenvalues be ordered such that
[…]
and
$\operatorname{Im}(\lambda_{m_{-}+1})\leq\operatorname{Im}(\lambda_{m_{-}+2})\leq\cdots\leq\operatorname{Im}(\lambda_{m_{-}+m_{0}}).$
Let $D_{i}(T)\in\mathbb{C}^{n_{i}\times n_{i}}$ be defined by
$D_{i}(T)\coloneq\left\{\begin{aligned} &I_{n_{i}}\qquad\text{if $i=1,\ldots,m_{-}$},\\ &\operatorname{diag}(T^{n_{i}-\frac{1}{2}},T^{n_{i}-\frac{3}{2}},\ldots,T^{\frac{1}{2}})\\ &\qquad\qquad\text{if $i=m_{-}+1,\ldots,m_{-}+m_{0}$},\\ &\mathrm{e}^{J_{i}T}\qquad\text{if $i=m_{-}+m_{0}+1,\ldots,m$},\end{aligned}\right$.
and define a scaling matrix $D(T)\in\mathbb{C}^{n\times n}$ by
$D(T)\coloneq\begin{pmatrix}D_{1}(T)\\ &\ddots\\ &&D_{m}(T)\end{pmatrix}$. (4)
We then define the scaled matrix
$W(B;T)\coloneq D(T)^{-1}Q^{-1}\widetilde{W}(B;T)Q^{-*}D(T)^{-*}.$ (5)
In what follows, we abbreviate $W(\operatorname{diag}(p);T)$ as $W(p;T)$ for $p\in\mathbb{R}^{n}$, and refer to it as the scaled controllability Gramian.
We further introduce additional notation for later use. Let $n_{-}\coloneq\sum_{i=1}^{m_{-}}n_{i},\ n_{0}\coloneq\sum_{i=m_{-}+1}^{m_{-}+m_{0}}n_{i}$, and $n_{+}\coloneq\sum_{i=m_{-}+m_{0}+1}^{m}n_{i}$ denote the numbers of eigenvalues with negative, zero, and positive real parts, respectively, counted with algebraic multiplicity. Let
$J_{-}\coloneq\begin{pmatrix}J_{1}\\ &\ddots\\ &&J_{m_{-}}\end{pmatrix}\in\mathbb{C}^{n_{-}\times n_{-}}$
be the Jordan matrix corresponding to the eigenvalues with negative real parts. Similarly, we define $J_{0}\in\mathbb{C}^{n_{0}\times n_{0}}$ and $J_{+}\in\mathbb{C}^{n_{+}\times n_{+}}$.
Furthermore, let $m^{\prime}_{0}$ denote the number of distinct values among $\lambda_{m_{-}+1},\ldots,\lambda_{m_{-}+m_{0}}$, and let $\lambda^{\prime}_{1},\cdots,\lambda^{\prime}_{m^{\prime}_{0}}$ be these values, ordered so that $\operatorname{Im}(\lambda^{\prime}_{1})<\cdots<\operatorname{Im}(\lambda^{\prime}_{m^{\prime}_{0}})$. We define the geometric multiplicity of $\lambda^{\prime}_{i}$ as $m_{0}^{(i)}\coloneq\sum_{\lambda_{j}=\lambda^{\prime}_{i}}1$, that is, the number of Jordan blocks associated with $\lambda_{i}^{\prime}$, and the algebraic multiplicity of $\lambda^{\prime}_{i}$ as $n_{0}^{(i)}\coloneq\sum_{\lambda_{j}=\lambda^{\prime}_{i}}n_{j}$.
III-B Infinite-horizon scaled controllability Gramian
We discuss the limit of the scaled controllability Gramian $W(B;T)$ as $T\to\infty$. We repeatedly use the following lemma.
Lemma 1 ([11, Theorem 2.4.4.1])
Let $C_{1}\in\mathbb{C}^{\ell_{1}\times\ell_{1}},C_{2}\in\mathbb{C}^{\ell_{2}\times\ell_{2}}$, and $P\in\mathbb{C}^{\ell_{1}\times\ell_{2}}$. If $\mu_{1}+\mu_{2}\neq 0$ holds for any eigenvalues $\mu_{1}$ of $C_{1}$ and $\mu_{2}$ of $C_{2}$, then
$C_{1}Y+YC_{2}+P=0$
has a unique solution $Y\in\mathbb{C}^{\ell_{1}\times\ell_{2}}$.
Let us define $\widetilde{B}\coloneq Q^{-1}BQ^{-*}$ and partition $\widetilde{B}$ into blocks in the following two ways:
$\widetilde{B}=\begin{pmatrix}\widetilde{B}_{1,1}&\cdots&\widetilde{B}_{1,m}\\ \vdots&\ddots&\vdots\\ \widetilde{B}_{m,1}&\cdots&\widetilde{B}_{m,m}\end{pmatrix}\\ =\begin{pmatrix}\widetilde{B}_{-,-}&\widetilde{B}_{-,0}&\widetilde{B}_{-,+}\\ \widetilde{B}_{0,-}&\widetilde{B}_{0,0}&\widetilde{B}_{0,+}\\ \widetilde{B}_{+,-}&\widetilde{B}_{+,0}&\widetilde{B}_{+,+}\end{pmatrix}$, (6)
where $\widetilde{B}_{i,j}\in\mathbb{C}^{n_{i}\times n_{j}}\ (i,j=1,\ldots,m)$, and $\widetilde{B}_{\alpha,\beta}\in\mathbb{C}^{n_{\alpha}\times n_{\beta}}\ (\alpha,\beta\in\{-,0,+\})$. Then, let $W_{-}(B)$ and $W_{+}(B)$ be the unique solutions to
$J_{-}W_{-}(B)+W_{-}(B)J_{-}^{*}+\widetilde{B}_{-,-}=0,$ (7)
$(-J_{+})W_{+}(B)+W_{+}(B)(-J_{+}^{*})+\widetilde{B}_{+,+}=0.$ (8)
For $i,j=m_{-}+1,\ldots,m_{-}+m_{0}$, we denote the $(n_{i},n_{j})$-entry of $\widetilde{B}_{i,j}$ by $\widetilde{b}_{i,j}$, which is the lower-right entry, and define
$C_{i,j}\coloneq\begin{pmatrix}\frac{1}{(n_{i}-1)!(n_{j}-1)!(n_{i}+n_{j}-1)}&\cdots&\frac{1}{(n_{i}-1)!0!n_{i}}\\ \vdots&\ddots&\vdots\\ \frac{1}{0!(n_{j}-1)!n_{j}}&\cdots&\frac{1}{0!0!1}\end{pmatrix}$. (9)
Moreover, define $W_{0}^{(\ell)}(B)\in\mathbb{C}^{n_{0}^{(\ell)}\times n_{0}^{(\ell)}}\ (\ell=1,\ldots,m_{0}^{\prime})$ and $W_{0}(B)\in\mathbb{C}^{n_{0}\times n_{0}}$ by
$W_{0}^{(\ell)}(B)\coloneq\begin{pmatrix}\widetilde{b}_{\alpha+1,\alpha+1}C_{\alpha+1,\alpha+1}&\cdots&\widetilde{b}_{\alpha+1,\beta}C_{\alpha+1,\beta}\\ \vdots&\ddots&\vdots\\ \widetilde{b}_{\beta,\alpha+1}C_{\beta,\alpha+1}&\cdots&\widetilde{b}_{\beta,\beta}C_{\beta,\beta}\end{pmatrix}$, (10)
$W_{0}(B)\coloneq\begin{pmatrix}W_{0}^{(1)}(B)\\ &\ddots\\ &&W_{0}^{(m_{0}^{\prime})}(B)\end{pmatrix}$, (11)
where we set $\alpha=m_{-}+m_{0}^{(1)}+\cdots+m_{0}^{(\ell-1)},\ \beta=m_{-}+m_{0}^{(1)}+\cdots+m_{0}^{(\ell)}$ to simplify the indexing.
Then, the limit of $W(B;T)$ can be expressed in terms of $W_{-}(B),W_{0}(B)$, and $W_{+}(B)$.
Theorem 1
The scaled controllability Gramian $W(B;T)$ converges as $T\to\infty$, and its limit is given by
$W_{\infty}(B)\coloneq\begin{pmatrix}W_{-}(B)\\ &W_{0}(B)\\ &&W_{+}(B)\end{pmatrix}$. (12)
Proof
See Appendix A. $\Box$
When the matrix $A$ has no eigenvalues on the imaginary axis, $W_{\infty}(B)$ coincides with the controllability Gramian defined in the frequency domain, up to a fixed transformation, which has been used in model reduction of unstable systems [31]. Since the infinite-horizon scaled controllability Gramian remains applicable even when $A$ has eigenvalues on the imaginary axis, it can be viewed as a generalization of the frequency-domain controllability Gramian. To the best of our knowledge, this work is the first to clarify the relationship between the limit $T\to\infty$ of controllability-Gramian-based optimization problems and these Gramians.
Analogously to $W(p;T)$, we abbreviate $W_{\infty}(\operatorname{diag}(p))$ as $W_{\infty}(p)$ for $p\in\mathbb{R}^{n}$, and refer to it as the infinite-horizon scaled controllability Gramian. The same notation is used for $W_{-},W_{0}$, and $W_{+}$.
For the finite-horizon controllability scoring problem, the following lemma guarantees that the feasible set is nonempty.
Lemma 2
Assume that $T>0$ is finite, and let $p\in\Delta$ and $p_{i}>0\ (i=1,\ldots,n)$. Then, $\widetilde{W}(p;T)\succ O$; equivalently, $W(p;T)\succ O$.
In the subsequent sections, we study controllability scoring problems based on matrices such as $W_{\infty}(p)$. The following lemma, analogous to Lemma 2, guarantees that the feasible set is nonempty.
Lemma 3
Let $p\in\Delta$ and $p_{i}>0\ (i=1,\ldots,n)$. Then, $W_{\infty}(p)\succ O$. In particular, $W_{-}(p)\succ O$.
Proof (Lemmas 2 and 3)
See Appendix B. $\Box$
III-C Reformulation of controllability scoring problems
We verify that Problems 3 and 4 are equivalent to Problems 1 and 2, respectively. It follows from 5 that
$\widetilde{W}(p;T)=QD(T)W(p;T)D(T)^{*}Q^{*}$
for $p\in\mathbb{R}^{n}$. Consequently,
$\log\det\widetilde{W}(p;T)\\ =-\log\det W(p;T)-\log\det\left(QD(T)D(T)^{*}Q^{*}\right),$
where the second term is independent of $p\in\mathbb{R}^{n}$. Moreover, since $Q$ and $D(T)$ are invertible, $\widetilde{W}(p;T)\succ O$ and $W(p;T)\succ O$ are equivalent. Therefore, Problem 3 is equivalent to Problem 1. Similarly, since
$\operatorname{tr}\left(\widetilde{W}(p;T)^{-1}\right)\\ =\operatorname{tr}\left(Q^{-*}D(T)^{-*}W(p;T)^{-1}D(T)^{-1}Q^{-1}\right),$
Problem 4 is equivalent to Problem 2.
Since the uniqueness of the optimal solutions is equivalent as well, we obtain the same result as in [23, Theorem 5] and [22, Theorem 1]. In Appendix E, we summarize existing results on the uniqueness of the optimal solutions and present new results.
If $A$ is not Hurwitz, $\widetilde{W}(p;T)$ diverges, and hence evaluating the objective functions of Problems 1 and 2 becomes numerically unstable. On the other hand, from Theorem 1, $W(p;T)$ converges to $W_{\infty}(p)$, and it is positive definite if $p_{i}>0\ (i=1,\ldots,n)$. Therefore, the objective function of Problem 3 remains numerically stable even for large $T$. Moreover, since $D(T)^{-1}$ converges, the objective function of Problem 4 is also numerically stable. Hence, ii) in Section II-B is thereby resolved.
IV Extension to infinite-horizon settings
In this section, we formalize the infinite-horizon controllability scoring problems by taking the limit $T\to\infty$ in Problems 3 and 4. We then justify this formulation by proving that, under suitable assumptions, the optimal solutions are uniquely determined and that the finite-horizon controllability scores indeed converge to them.
In what follows, we define
$W_{\infty,i}\coloneq W_{\infty}(e_{i}e_{i}^{\top}),$
for which
$W_{\infty}(p)=\sum_{i=1}^{n}p_{i}W_{\infty,i}$
holds for $p=(p_{1},\ldots,p_{n})^{\top}\in\mathbb{R}^{n}$. We define $W_{-,i},W_{0,i}$, and $W_{+,i}$ analogously.
IV-A Problem formulation
We formulate the infinite-horizon version of Problem 3 as follows.
Problem 5
$\operatorname*{minimize}_{p}\quad$ ${-}\log\det W_{\infty}(p)$
$\mathrm{subject\ to}\quad$ $p\in\Delta,\ W_{\infty}(p)\succ O.$
Since the infinite-horizon version of Problem 4 is not well defined unless eigenvalues with negative real parts are present, we impose the following assumption.
Assumption 1
The matrix $A$ has at least one eigenvalue with a negative real part, i.e., $n_{-}\geq 1$.
Under Assumption 1, we formulate the infinite-horizon version of Problem 4 as follows.
Problem 6
$\operatorname*{minimize}_{p}\quad$ $\operatorname{tr}\left(Q^{-*}\begin{pmatrix}W_{-}(p)^{-1}\\ &O\end{pmatrix}Q^{-1}\right)$
$\mathrm{subject\ to}\quad$ $p\in\Delta,\ W_{-}(p)\succ O.$
We refer to the optimal solution to Problem 5 as the VCS on an infinite time horizon, or simply the infinite-horizon VCS. Similarly, we refer to the optimal solution to Problem 6 as the AECS on an infinite time horizon, or simply the infinite-horizon AECS. In the case where $A$ is Hurwitz, Problems 5 and 6 coincide with the controllability scoring problems proposed in [23].
Since $W(p;T)\to W_{\infty}(p)$, Problem 5 is formally obtained from Problem 3 by letting $T\to\infty$, and it depends on the whole matrix $W_{\infty}(p)$. On the other hand, since $D(T)^{-1}\to\begin{pmatrix}I_{n_{-}}\\ &O\end{pmatrix}$, formally letting $T\to\infty$ in Problem 4 yields the objective function of Problem 6, which depends only on $W_{-}(p)$; consequently, the constraint is posed only on $W_{-}(p)$.
Remark 2
Problem 5 imposes a constraint that ensures controllability of the entire system, whereas Problem 6 requires controllability only of the stable modes. Consequently, while the infinite-horizon VCS enforces controllability of the full system, the infinite-horizon AECS may render the entire system uncontrollable.
IV-B Convergence of controllability scores
Problems 5 and 6 are obtained by formally taking the limit $T\to\infty$ in Problems 1 and 2. Under suitable assumptions, the optimal solutions to Problems 1 and 2 (finite-horizon controllability scores) indeed converge to those to Problems 5 and 6 (infinite-horizon controllability scores).
Theorem 2
Assume that Problem 5 admits a unique optimal solution $p_{\mathrm{VCS}}^{\infty}$. Then, the finite-horizon controllability score $p_{\mathrm{VCS}}^{T}$, which is the optimal solution to Problem 1, converges to $p_{\mathrm{VCS}}^{\infty}$ as $T\to\infty$.
Theorem 3
Assume that Assumption 1 holds, and that Problem 6 admits a unique optimal solution $p_{\mathrm{AECS}}^{\infty}$. Then, the finite-horizon controllability score $p_{\mathrm{AECS}}^{T}$, which is the optimal solution to Problem 2, converges to $p_{\mathrm{AECS}}^{\infty}$ as $T\to\infty$.
Proof (Theorems 2 and 3)
See Appendix C. $\Box$
Theorems 2 and 3 assume the uniquenss of the optimal solutions to Problems 5 and 6. As discussed in Section IV-C, this assumption is satisfied under appropriate conditions.
IV-C Uniqueness analysis
We discuss the uniqueness of the optimal solutions to Problems 5 and 6. The uniqueness of the optimal solution is important for two reasons.
First, the convergence results for the finite-horizon controllability scores established in Section IV-B are derived under the assumption that the optimal solutions to Problems 5 and 6 are unique. When the optimal solutions are unique, the finite-horizon controllability scores converge to them as $T\to\infty$; however, if the optimal solutions are not unique, their asymptotic behavior as $T\to\infty$ is generally unclear.
Second, uniqueness is essential from the viewpoints of reproducibility and interpretability. As has been studied in detail for finite-horizon controllability scores [23, 22], the uniqueness is crucial when they are used as centrality measures. Existing results on the uniqueness of finite-horizon controllability scores are summarized in Appendix E. If the optimal solution is not unique, different researchers may obtain different infinite-horizon controllability scores and consequently draw different conclusions, which undermines reproducibility. Moreover, the presence of multiple optimal solutions raises interpretability issues regarding how such solutions should be used to define a meaningful centrality measure.
We show that although the uniqueness of the optimal solution depends on the structure of the eigenvectors of the matrix $A$, it holds in a wide range of practically relevant cases. Note that even if the finite-horizon controllability score is unique for all $T>0$, the infinite-horizon controllability score need not be unique. Indeed, the following example shows that the finite-horizon controllability scores are unique for all $T>0$, whereas the infinite-horizon controllability scores are not unique.
Example 1
Let $A=\begin{pmatrix}0&1\\ 1&0\end{pmatrix}$. Then, Problems 1 and 2 each admit a unique optimal solution for all $T>0$. However, Problems 5 and 6 have multiple optimal solutions.
Indeed, by taking $Q=\dfrac{1}{\sqrt{2}}\begin{pmatrix}1&-1\\ 1&1\end{pmatrix}$, we obtain $Q^{-1}AQ=\operatorname{diag}(1,-1)$, whose eigenvalues are $1$ and 1-1. Therefore, by Theorem 7, Problems 1 and 2 each admit a unique optimal solution for all $T>0$.
On the other hand, the infinite-horizon scaled controllability Gramians, are given by $W_{\infty,1}=W_{\infty,2}=\begin{pmatrix}\frac{1}{4}&0\\ 0&\frac{1}{4}\end{pmatrix}$ and $W_{-,1}=W_{-,2}=\frac{1}{4}$. Thus, $W_{\infty}(p)$ and $W_{-}(p)$ take constant matrices for all $p\in\Delta$. As a consequence, every point in $\Delta$ is an optimal solution to Problems 5 and 6. $\Box$
We partition $Q^{-1}e_{i}$ into blocks in following two ways:
$Q^{-1}e_{i}=\begin{pmatrix}q_{1,i}\\ \vdots\\ q_{m,i}\end{pmatrix}=\begin{pmatrix}q_{-,i}\\ q_{0,i}\\ q_{+,i}\end{pmatrix}$.
Furthermore, let $q^{\prime}_{k,i}$ denote the $n_{k}$-th (i.e., the last) entry of $q_{k,i}$. For $\ell=1,\ldots,m_{0}^{\prime}$, we define
$q_{0,i}^{(\ell)}\coloneq(q^{\prime}_{\alpha+1,i},\ldots,q^{\prime}_{\beta,i})^{\top}\in\mathbb{C}^{m_{0}^{(\ell)}},$
where we set $\alpha=m_{-}+m_{0}^{(1)}+\cdots+m_{0}^{(\ell-1)},\ \beta=m_{-}+m_{0}^{(1)}+\cdots+m_{0}^{(\ell)}$ to simplify the indexing. Then, the following theorems hold.
Theorem 4
Let
$M_{\mathrm{VCS}}$ $\coloneq\begin{pmatrix}q_{-,1}\otimes q_{-,1}&\cdots&q_{-,n}\otimes q_{-,n}\\ q_{0,1}^{(1)}\otimes q_{0,1}^{(1)}&\cdots&q_{0,n}^{(1)}\otimes q_{0,n}^{(1)}\\ \vdots&\ddots&\vdots\\ q_{0,1}^{(m_{0}^{\prime})}\otimes q_{0,1}^{(m_{0}^{\prime})}&\cdots&q_{0,n}^{(m_{0}^{\prime})}\otimes q_{0,n}^{(m_{0}^{\prime})}\\ q_{+,1}\otimes q_{+,1}&\cdots&q_{+,n}\otimes q_{+,n}\end{pmatrix}$
$\in\mathbb{C}^{(n_{-}^{2}+(m_{0}^{(1)})^{2}+\cdots+(m_{0}^{(m_{0}^{\prime})})^{2}+n_{+}^{2})\times n},$
and assume $\operatorname{rank}M_{\mathrm{VCS}}=n$. Then, Problem 5 admits a unique optimal solution.
Theorem 5
Let
$M_{\mathrm{AECS}}\coloneq\begin{pmatrix}q_{-,1}\otimes q_{-,1}&\cdots&q_{-,n}\otimes q_{-,n}\end{pmatrix}\in\mathbb{C}^{n_{-}^{2}\times n},$
and assume $\operatorname{rank}M_{\mathrm{AECS}}=n$. Then, Problem 6 admits a unique optimal solution.
Proof (Theorems 4 and 5)
See Appendix D. $\Box$
In many applications, the number of unstable modes of $A$ is small, so that $n_{-}\approx n$, and consequently the numbers of rows of $M_{\mathrm{VCS}}$ and $M_{\mathrm{AECS}}$ become much larger than the numbers of columns of them. If $M_{\mathrm{VCS}}$ and $M_{\mathrm{AECS}}$ have full column rank, then the optimal solutions to Problems 5 and 6 are unique. Although the structure of the eigenvectors depends on the matrix $A$, if the eigenvectors are drawn generically from $\mathbb{C}^{n}$, $M_{\mathrm{VCS}}$ and $M_{\mathrm{AECS}}$ are expected to have full column rank in almost all cases.
V Algorithms
In this section, we describe an algorithm for solving Problems 5 and 6. In formulating Problems 5 and 6, we employed the Jordan canonical form; however, a direct use of the Jordan canonical form may lead to numerical instability [8]. As shown below, under the following assumption, the optimal solutions to Problems 5 and 6 can be computed in a numerically stable manner via the real Schur decomposition and the solutions of Sylvester equations.
Assumption 2
The only eigenvalue of $A$ on the imaginary axis is zero, and it is semisimple; that is, $A$ has no purely imaginary eigenvalues, and if $A$ has zero as an eigenvalue, then all associated Jordan blocks are of size one.
An important example satisfying this assumption is the graph Laplacian.
Example 2
The matrix $L=(\ell_{ij})\in\mathbb{R}^{n\times n}$ is said to be the graph Laplacian if, for all $i\neq j$,
[…]
and the diagonal entries satisfy
$\ell_{ii}=-\sum_{j\neq i}\ell_{ij}.$
When considering graph Laplacians, we restrict attention to positive edge weights whenever an edge exists. For undirected graphs, the Laplacian matrix $L$ is symmetric, whereas for directed graphs, $L$ is generally not symmetric.
Let $L$ be the Laplacian matrix of a directed graph. Then, A=LA=-L satisfies Assumption 2. In fact, from the definition, the matrix $A=(a_{ij})$ satisfies
$a_{ij}\geq 0\ (i\neq j),\quad a_{ii}=-\sum_{j\neq i}a_{ij}\leq 0.$
It follows from Geršhgorin’s theorem [27, Theorem 1.1] that all eigenvalues of $A$ have nonpositive real parts, and that zero is the only eigenvalue that may lie on the imaginary axis. Furthermore, $\operatorname{rank}(A)=\operatorname{rank}(A^{2})$ holds [6, Proposition 12]. This implies that the zero eigenvalue of $A$ is semisimple. $\Box$
First, we perform a block diagonalization using the real Schur decomposition and the solution of a Sylvester equation. Since this approach does not require the Jordan canonical form, it is numerically stable. By the real Schur decomposition, we can find an orthogonal matrix $R^{(1)}\in\mathbb{R}^{n\times n}$ such that
$\left(R^{(1)}\right)^{\top}AR^{(1)}=\begin{pmatrix}A_{-}\\ A_{0,-}&A_{0}\\ A_{+,-}&A_{+,0}&A_{+}\end{pmatrix}$, (13)
where $A_{-},A_{0}$, and $A_{+}$ have eigenvalues with negative, zero, and positive real parts, respectively. While the real Schur decomposition typically yields a block upper-triangular matrix, applying it to $A^{\top}$ yields a block lower-triangular matrix.
Let $R^{(2)}_{0,-},R^{(2)}_{+,-}$, and $R^{(2)}_{+,0}$ be matrices satisfying
$A_{0}R^{(2)}_{0,-}-R^{(2)}_{0,-}A_{-}=-A_{0,-},$ (14)
$A_{+}R^{(2)}_{+,0}-R^{(2)}_{+,0}A_{0}=-A_{+,0},$ (15)
$A_{+}R^{(2)}_{+,-}-R^{(2)}_{+,-}A_{-}=-(A_{+,-}+A_{+,0}R^{(2)}_{0,-})$. (16)
Since no two of the matrices $A_{-},A_{0}$, and $A_{+}$ share a common eigenvalue, 14, 15, and 16 can be solved successively and admit a unique solution from Lemma 1. Define
$R^{(2)}\coloneq\begin{pmatrix}I_{n_{-}}\\ R^{(2)}_{0,-}&I_{n_{0}}\\ R^{(2)}_{+,-}&R^{(2)}_{+,0}&I_{n_{+}}\end{pmatrix}$
and set $R=R^{(1)}R^{(2)}$. Then, by using
$\left(R^{(2)}\right)^{-1}=\begin{pmatrix}I_{n_{-}}\\ -R^{(2)}_{0,-}&I_{n_{0}}\\ -R^{(2)}_{+,-}+R^{(2)}_{+,0}R^{(2)}_{0,-}&-R^{(2)}_{+,0}&I_{n_{+}}\end{pmatrix}$, (17)
we obtain the block-diagonal form
$R^{-1}AR=\begin{pmatrix}A_{-}\\ &A_{0}\\ &&A_{+}\end{pmatrix}$. (18)
Then, let
$\begin{pmatrix}r_{-,i}\\ r_{0,i}\\ r_{+,i}\end{pmatrix}\coloneq R^{-1}e_{i},$ (19)
where $r_{-,i}\in\mathbb{R}^{n_{-}},\ r_{0,i}\in\mathbb{R}^{n_{0}},\ r_{+,i}\in\mathbb{R}^{n_{+}}$, and let $V_{-,i}\in\mathbb{R}^{n_{-}\times n_{-}}$ and $V_{+,i}\in\mathbb{R}^{n_{+}\times n_{+}}$ be the solutions to
$A_{-}V_{-,i}+V_{-,i}A_{-}^{\top}+q_{-,i}r_{-,i}^{\top}=0,$ (20)
$A_{+}V_{+,i}+V_{+,i}A_{+}^{\top}-q_{+,i}r_{+,i}^{\top}=0.$ (21)
Moreover, let $V_{0,i}\coloneq r_{0,i}r_{0,i}^{\top}$. Finally, for $p=(p_{1},\ldots,p_{n})^{\top}\in\mathbb{R}^{n}$, we define
$V_{\alpha}(p)\coloneq\sum_{i=1}^{n}p_{i}V_{\alpha,i}\quad(\alpha\in\{-,0,+\}),$
$V_{i}\coloneq\begin{pmatrix}V_{-,i}\\ &V_{0,i}\\ &&V_{+,i}\end{pmatrix}$,
$V(p)\coloneq\begin{pmatrix}V_{-}(p)\\ &V_{0}(p)\\ &&V_{+}(p)\end{pmatrix}$.
We then consider the following alternative problems.
Problem 7
$\operatorname*{minimize}_{p}\quad$ ${-}\log\det V(p)$
$\mathrm{subject\ to}\quad$ $p\in\Delta,\ V(p)\succ O.$
Problem 8
$\operatorname*{minimize}_{p}\quad$ $\operatorname{tr}\left(V_{-}(p)^{-1}\right)$
$\mathrm{subject\ to}\quad$ $p\in\Delta,\ V_{-}(p)\succ O.$
In fact, Problems 5 and 6 are equivalent to Problems 7 and 8, respectively.
Theorem 6
Under Assumption 2, Problems 5 and 7 are equivalent in the sense that they share the same set of optimal solutions. Under Assumptions 1 and 2, Problems 6 and 8 are also equivalent.
Proof
Since $A$ is block-diagonalized as in 18, the Jordan canonical forms of $A_{-},A_{0}$, and $A_{+}$ are given by $J_{,}J_{0}$, and $J_{+}$, respectively. Thus, there exist invertible matrices $S_{-}\in\mathbb{C}^{n_{-}\times n_{-}},\ S_{0}\in\mathbb{C}^{n_{0}\times n_{0}}$, and $S_{+}\in\mathbb{C}^{n_{+}\times n_{+}}$ such that $S_{-}^{-1}A_{-}S_{-}=J_{-},\ S_{0}^{-1}A_{0}S_{0}=J_{0}$, and $S_{+}^{-1}A_{+}S_{+}=J_{+}$. Moreover, letting
$S\coloneq\begin{pmatrix}S_{-}\\ &S_{0}\\ &&S_{+}\end{pmatrix}$,
we obtain $S^{-1}(R^{-1}AR)S=J$; hence $S_{-},S_{0}$ and $S_{+}$ can be chosen so that $RS=Q$.
As defined in Section IV-C,
$\begin{pmatrix}q_{-,i}\\ q_{0,i}\\ q_{+,i}\end{pmatrix}=Q^{-1}e_{i}=S^{-1}\begin{pmatrix}r_{-,i}\\ r_{0,i}\\ r_{+,i}\end{pmatrix}$.
Then, $r_{-,i}=S_{-}q_{-,i},\ r_{0,i}=S_{0}q_{0,i}$, and $r_{+,i}=S_{+}q_{+,i}$ hold.
From 7 and 8, $W_{-,i}$ and $W_{+,i}$ are the unique solutions to
$J_{-}W_{-,i}+W_{-,i}J_{-}^{*}+r_{-,i}r_{-,i}^{*}=0,$ (22)
$(-J_{+})W_{+,i}+W_{+,i}(-J_{+}^{*})+r_{+,i}r_{+,i}^{*}=0,$ (23)
respectively. Multiplying 22 from the left by $S_{-}$ and from the right by $S_{-}^{*}$, and using the uniqueness of solutions to 20, we obtain $V_{-,i}=S_{-}W_{-,i}S_{-}^{*}$. Similary, it follows from 21 and 23 that $V_{+,i}=S_{+}W_{+,i}S_{+}^{*}$. Since $W_{0,i}$ reduces to $q_{0,i}q_{0,i}^{*}$ under Assumption 2, we obtain $V_{0,i}=S_{0}W_{0,i}S_{0}^{*}$. Therefore,
$W_{\infty}(p)=S^{-1}V(p)S^{-*},$ (24)
$\begin{pmatrix}W_{-}(p)^{-1}\\ &O\end{pmatrix}=S^{*}\begin{pmatrix}V_{-}(p)^{-1}\\ &O\end{pmatrix}S$ (25)
hold.
For the objective functions of Problems 5 and 7, since
$\log\det W_{\infty}(p)=-\log\det V(p)-\log\det\left(S^{-1}S^{-*}\right),$
their difference is a constant independent of $p$. Moreover, since $W_{\infty}(p)\succ O$ is equivalent to $V(p)\succ O$, Problems 5 and 7 are equivalent.
By using $SQ^{-1}=R^{-1}=(R^{(2)})^{-1}(R^{(1)})^{\top}$ and 17, we obtain
$\operatorname{tr}\left(Q^{-*}\begin{pmatrix}W_{-}(p)^{-1}\\ &O\end{pmatrix}Q^{-1}\right)$
$=\operatorname{tr}\left(Q^{-*}S^{*}\begin{pmatrix}V_{-}(p)^{-1}\\ &O\end{pmatrix}SQ^{-1}\right)$
$=\operatorname{tr}\left(R^{(1)}(R^{(2)})^{-\top}\begin{pmatrix}V_{-}(p)^{-1}\\ &O\end{pmatrix}(R^{(2)})^{-1}(R^{(1)})^{\top}\right)$
$=\operatorname{tr}\left(R^{(1)}\begin{pmatrix}V_{-}(p)^{-1}\\ &O\end{pmatrix}(R^{(1)})^{\top}\right)$
$=\operatorname{tr}\left(V_{-}(p)^{-1}\right),$
where we used the fact that $R^{(1)}$ is orthogonal and that $R^{(2)}$ is lower triangular. Consequently, the objective functions of Problems 6 and 8 are equal. Moreover, $W_{-}(p)\succ O$ is equivalent to $V_{-}(p)\succ O$. Hence, Problems 6 and 8 are equivalent. $\Box$
Although Problems 5 and 6 are formulated based on the Jordan canonical form, under Assumptions 1 and 2, they are respectively equivalent to Problems 7 and 8. Hence, it suffices to compute the optimal solutions to Problems 7 and 8, without resorting to the Jordan canonical form.
To summarize the discussion so far, under Assumptions 1 and 2, $V_{-,i},V_{0,i}$, and $V_{+,i}$ can be computed by Algorithm 1.
0: The matrix $A$.
1: Perform the real Schur decomposition 13 and compute $R^{(1)}$.
2: Compute $R^{(2)}_{0,-},R^{(2)}_{+,-}$, and $R^{(2)}_{+,0}$ by 14, 15, and 16, and then compute $(R^{(2)})^{-1}$ by 17.
3: Compute $r_{-,i},r_{0,i}$, and $r_{+,i}$ by 19.
4: Compute $V_{-,i}$ by solving 20.
[…]
6: Compute $V_{+,i}$ by solving 21.
[…]
Algorithm 1 The computation of $V_{-,i},V_{0,i}$, and $V_{+,i}$
Once $V_{\mathrm{l},i},V_{\mathrm{i},i}$, and $V_{\mathrm{r},i}$ have been computed, Problems 7 and 8 take the same form as the Problems 1 and 2. Hence, we can employ Algorithms 2 and 3, which are the same as those in [23]. Here, $f$ is the objective function, and $\Pi_{\Delta}$ is a Euclidean projection onto $\Delta$.
0: The terminal condition $\varepsilon\geq 0$ and the initial point $p^{(0)}\coloneq(1/n,\ldots,1/n)^{\top}$.
1: Compute $V_{-,i},V_{0,i},V_{+,i}\ (i=1,\ldots,n)$ by Algorithm 1.
2:for $k=0,1,\ldots$ do
3:  Determine a step size $\alpha^{(k)}$ by Algorithm 3.
4:  $p^{(k+1)}\leftarrow\Pi_{\Delta}\left(p^{(k)}-\alpha^{(k)}\nabla f(p^{(k)})\right)$
5:  if $\left\|{p^{(k)}-p^{(k+1)}}\right\|\leq\varepsilon$ then
6:   return $p^{(k+1)}$
7:  end if
8:end for
Algorithm 2 The projected gradient method for Problems 7 and 8
0: The point $p^{(k)}$ in Algorithm 2, the parameters $\sigma,\rho\in(0,1)$, and the initial step size $\alpha>0$.
[…]
2:while true do
3:  $p^{\prime(k)}\leftarrow\Pi_{\Delta}\left(p^{(k)}-\alpha^{(k)}\nabla f(p^{(k)})\right)$
4:  if $f(p^{\prime(k)})\leq f(p^{(k)})+\sigma\nabla f(p^{(k)})^{\top}(p^{\prime(k)}-p^{(k)})$ then
5:   break
6:  else
7:   $\alpha^{(k)}\leftarrow\rho\alpha^{(k)}$
8:  end if
9:end while
10:return $\alpha^{(k)}$
Algorithm 3 The Armijo rule
Let the objective functions of Problems 7 and 8 be
$g(p)\coloneq-\log\det V(p),$
$h(p)\coloneq\operatorname{tr}\left(V_{-}(p)^{-1}\right).$
Then, their gradients can be computed by
$(\nabla g(p))_{i}=\operatorname{tr}\left(V(p)^{-1}V_{i}\right),$
$(\nabla h(p))_{i}=\operatorname{tr}(V_{-}(p)^{-1}V_{-,i}V_{-}(p)^{-1}).$
The feasible sets for Problems 7 and 8 are given by
$\mathcal{F}_{\mathrm{VCS}}\coloneq\{p\in\mathbb{R}^{n}\mid p\in\Delta,\ V(p)\succ O\},$
$\mathcal{F}_{\mathrm{AECS}}\coloneq\{p\in\mathbb{R}^{n}\mid p\in\Delta,\ V_{-}(p)\succ O\},$
respectively. It follows from Lemmas 3 and 6 that $p^{(0)}=(1/n,\ldots,1/n)^{\top}\in\mathcal{F}_{\mathrm{VCS}}$ and $p^{(0)}\in\mathcal{F}_{\mathrm{AECS}}$. In the VCS case, once $V(p^{(0)})\succ O$ holds, the generated sequence $\{p^{(k)}\}_{k\in\mathbb{N}}$ satisfies $V(p^{(k)})\succ O$ inductively, and hence $p^{(k)}\in\mathcal{F}_{\mathrm{VCS}}$, since $g(p^{(k+1)})\leq g(p^{(k)})$ holds [3, Proposition 2.1.3],[23]. A key point is that, instead of projecting onto $\mathcal{F}_{\mathrm{VCS}}$, we project onto $\Delta$, which can be computed efficiently [7]. Since the objective function is convex, convergence to the optimal solution is guaranteed [12, 23]. The same argument applies to the AECS case.
VI Numerical experiments
In this section, we conduct numerical experiments using the same network as in [22], which is illustrated in Fig. 2. The network consists of $n=10$ nodes, and all edges have a uniform weight $c=0.2$. Let $L$ denote the graph Laplacian of this network, and set A=LA=-L.
12345678910
Figure 2: Directed network employed in the numerical experiments
While [22] computes the controllability scores for finite horizons up to $T=10000$, we compute those on an infinite horizon using Problems 7 and 8. The results are compared in Table I.
TABLE I: Comparison of the controllability scores on finite horizons with those on an infinite horizon
(a) VCS
Node $T=0.01$ $T=1$ $T=1000$ $T=10000$ $T=\infty$
1 0.10000 0.09968 0.07335 0.07327 0.07329
2 0.10000 0.10000 0.10112 0.10108 0.10104
3 0.10000 0.10000 0.10876 0.10874 0.10871
4 0.10000 0.10000 0.08638 0.08636 0.08634
5 0.10000 0.09967 0.04556 0.04499 0.04490
6 0.10000 0.09935 0.06074 0.06071 0.06091
7 0.10000 0.10131 0.24929 0.24952 0.24967
8 0.10000 0.09967 0.04231 0.04221 0.04220
9 0.10000 0.10033 0.16614 0.16674 0.16686
10 0.10000 0.10000 0.06636 0.06632 0.06631
(b) AECS
Node $T=0.01$ $T=1$ $T=1000$ $T=10000$ $T=\infty$
1 0.10000 0.10927 0.17127 0.17281 0.17428
2 0.10000 0.10000 0.11333 0.11364 0.11408
3 0.10000 0.10000 0.12054 0.12093 0.12174
4 0.10000 0.10000 0.10584 0.10610 0.10610
5 0.10000 0.09978 0.09075 0.09230 0.09275
6 0.10000 0.10905 0.13350 0.13383 0.13433
7 0.10000 0.09128 0.09257 0.09275 0.09233
8 0.10000 0.09982 0.06947 0.06945 0.06901
9 0.10000 0.09982 0.00703 0.00234 0.00000
10 0.10000 0.09998 0.09571 0.09586 0.09538
In the case of the VCS, the difference between the values at $T=10000$ and at $T=\infty$ is approximately $10^{-4}$ for each state node. Consistent with Theorems 2 and 6, the results suggest that the finite-horizon VCS converges to the infinite-horizon VCS as $T\to\infty$.
For the AECS, the difference between the values at $T=10000$ and at $T=\infty$ is somewhat larger, being approximately $10^{-3}$. However, consistent with Theorems 3 and 6, the results also suggest convergence of the finite-horizon AECS to the infinite-horizon AECS as $T\to\infty$.
An interesting observation is that, in the infinite-horizon AECS, node 9 is assigned a value of zero. Since node 9 is not influenced by any other nodes, setting $p_{9}=0$ renders the overall system uncontrollable. Nevertheless, the infinite-horizon AECS assigns zero weight to node 9. As discussed in Remark 2, this phenomenon can be explained by the fact that the infinite-horizon AECS focuses only on the controllability of the stable modes. In contrast, the infinite-horizon VCS enforces controllability of the entire system.
As in the finite-horizon case [22], the infinite-horizon VCS and AECS exhibit substantially different behaviors. This difference can be partially explained by the difference in the imposed controllability constraints discussed above: while the VCS enforces controllability of the entire system, the AECS requires controllability only of the stable modes. A complete characterization of the discrepancy, however, requires further analysis and is left for future work.
VII Conclusion
VII-A Summary
In this paper, we introduced a scaled controllability Gramian and used it to reformulate conventional controllability scoring problems into forms that can be solved in a numerically stable manner even when the system matrix $A$ is non-Hurwitz. By further taking the limit $T\to\infty$ in these reformulated problems, we formulated infinite-horizon controllability scoring problems. Under suitable assumptions, we proved that the finite-horizon controllability scores converge to the infinite-horizon ones and that the infinite-horizon controllability scores are unique. This proves a principled choice of the terminal time, namely $T=\infty$.
Although the formulations of infinite-horizon controllability scoring problems rely on the Jordan canonical form of $A$, we also showed that, under appropriate assumptions, their optimal solutions can be computed using the Schur decomposition and the solution of Sylvester equations, and we presented a corresponding numerical algorithm. In the numerical experiments, we computed infinite-horizon controllability scores for the same Laplacian dynamics as in [22] and observed that finite-horizon controllability scores indeed converge to infinite-horizon ones as $T\to\infty$.
VII-B Discrepancy between VCS and AECS
We showed that the infinite-horizon VCS and AECS exhibit markedly different behaviors, both theoretically and numerically. In the numerical experiments, the VCS and AECS take substantially different values, consistent with the observations reported in [22]. This discrepancy can be partially attributed to the difference in the controllability constraints imposed by the two measures. While the infinite-horizon VCS enforces controllability of the entire system, the infinite-horizon AECS requires controllability only of the stable modes. As a consequence, the AECS may assign values that render the overall system uncontrollable, which is indeed observed in the numerical experiments. Although this difference in constraints provides a partial explanation for the distinct behaviors of the VCS and AECS, a complete understanding of their discrepancy requires further analysis and is left for future work.
VII-C Application to real data
In [22, 24], the controllability score was applied to brain networks, where it was reported that the VCS assigns higher importance to sensory and emotional regions, whereas the AECS emphasizes regions associated with cognitive and motor functions. While the controllability score is expected to provide insights into the functional organization of real-world networks, its applications remain limited. Applying the controllability score to a broader range of large-scale networks and analyzing their functional roles from a controllability perspective are left for future work.
Appendix A Proof of Theorem 1
To prove Theorem 1, we prepare the following lemma.
Lemma 4 ([14, Theorem 5])
Let $C_{1}\in\mathbb{C}^{\ell_{1}\times\ell_{1}},C_{2}\in\mathbb{C}^{\ell_{2}\times\ell_{2}}$, and $R\in\mathbb{C}^{\ell_{1}\times\ell_{2}}$. If $\operatorname{Re}(\mu_{1}+\mu_{2})<0$ holds for any eigenvalues $\mu_{1}$ of $C_{1}$ and $\mu_{2}$ of $C_{2}$, then
$\int_{0}^{\infty}\mathrm{e}^{C_{1}t}P\mathrm{e}^{C_{2}t}\mathrm{d}t$
converges, and it is the unique solution $Y\in\mathbb{C}^{\ell_{1}\times\ell_{2}}$ satisfying
$C_{1}Y+YC_{2}+P=0.$
Proof (Theorem 1)
We prove that $W(B;T)$ converges to $W_{\infty}(B)$ as $T\to\infty$. By definition of $\widetilde{B}$, we obtain
$\widetilde{W}(B;T)=Q\left(\int_{0}^{T}\mathrm{e}^{Jt}\widetilde{B}\mathrm{e}^{J^{*}t}\mathrm{d}t\right)Q^{*}.$ (26)
Then, using the block decomposition in 6, we obtain
$\mathrm{e}^{Jt}\widetilde{B}\mathrm{e}^{J^{*}t}=\begin{pmatrix}\mathrm{e}^{J_{1}t}\widetilde{B}_{1,1}\mathrm{e}^{J_{1}^{*}t}&\cdots&\mathrm{e}^{J_{1}t}\widetilde{B}_{1,m}\mathrm{e}^{J_{m}^{*}t}\\ \vdots&\ddots&\vdots\\ \mathrm{e}^{J_{m}t}\widetilde{B}_{m,1}\mathrm{e}^{J_{1}^{*}t}&\cdots&\mathrm{e}^{J_{m}t}\widetilde{B}_{m,m}\mathrm{e}^{J_{m}^{*}t}\end{pmatrix}$ (27)
From 4, 5, 26, and 27, the $(i,j)$-block of $W(B;T)$ can be written as
$D_{i}(T)^{-1}\left\{\int_{0}^{T}\mathrm{e}^{J_{i}t}\widetilde{B}_{i,j}\mathrm{e}^{J_{j}^{*}t}\mathrm{d}t\right\}D_{j}(T)^{-*}.$ (28)
In the following, we analyze the block representation 28 by distinguishing the following nine cases:
  1. 1)
    $i=1,\ldots,m_{-},\quad j=1,\ldots,m_{-}$
  2. 2)
    $i=1,\ldots,m_{-},\quad j=m_{-}+1,\ldots,m_{-}+m_{0}$
  3. 3)
    $i=1,\ldots,m_{-},\quad j=m_{-}+m_{0}+1,\ldots,m$
  4. 4)
    $i=m_{-}+1,\ldots,m_{-}+m_{0},\quad j=1,\ldots,m_{-}$
  5. 5)
    $i=m_{-}+1,\ldots,m_{-}+m_{0},\quad j=m_{-}+1,\ldots,m_{-}+m_{0}$
  6. 6)
    $i=m_{-}+1,\ldots,m_{-}+m_{0},\quad j=m_{-}+m_{0}+1,\ldots,m$
  7. 7)
    $i=m_{-}+m_{0}+1,\ldots,m,\quad j=1,\ldots,m_{-}$
  8. 8)
    $i=m_{-}+m_{0}+1,\ldots,m,\quad j=m_{-}+1,\ldots,m_{-}+m_{0}$
  9. 9)
    $i=m_{-}+m_{0}+1,\ldots,m,\quad j=m_{-}+m_{0}+1,\ldots,m$
Case 1)
Noting that $D_{i}(T)=I_{n_{i}}$ and $D_{j}(T)=I_{n_{j}}$, 28 reduces to
$\int_{0}^{T}\mathrm{e}^{J_{i}t}\widetilde{B}_{i,j}\mathrm{e}^{J_{j}^{*}t}\mathrm{d}t.$
Since $\operatorname{Re}(\lambda_{i}),\operatorname{Re}(\lambda_{j})<0$, it follows from Lemma 4 that the $(i,j)$-block converges, as $T\to\infty$, to the unique solution $W_{-}^{(i,j)}(B)\in\mathbb{C}^{n_{i}\times n_{j}}$ satisfying
$J_{i}W_{-}^{(i,j)}(B)+W_{-}^{(i,j)}(B)J_{j}^{*}+\widetilde{B}_{i,j}=0.$
Then, the upper-left block of $W(B;T)$ of size $n_{-}\times n_{-}$ converges to
$\begin{pmatrix}W_{-}^{(1,1)}(B)&\cdots&W_{-}^{(1,m_{-})}B)\\ \vdots&\ddots&\vdots\\ W_{-}^{(m_{-},1)}(B)&\cdots&W_{-}^{(m_{-},m_{-})}(B)\end{pmatrix}$,
which coincides with $W_{-}(B)$.
Case 2)
Noting that $D_{i}(T)=I_{n_{i}}$, 28 reduces to
$\left\{\int_{0}^{T}\mathrm{e}^{J_{i}t}\widetilde{B}_{i,j}\mathrm{e}^{J_{j}^{*}t}\mathrm{d}t\right\}D_{j}(T)^{-*}.$
Since $\operatorname{Re}(\lambda_{i})<0=\operatorname{Re}(\lambda_{j})$, it follows from Lemma 4 that the integral converges as $T\to\infty$. Moreover, since $D_{j}(T)^{-*}\to 0$, the $(i,j)$-block converges to zero.
Case 3)
Noting that $D_{i}(T)=I_{n_{i}}$ and $D_{j}(T)=\mathrm{e}^{J_{j}T}$, 28 reduces to
$\left\{\int_{0}^{T}\mathrm{e}^{J_{i}t}\widetilde{B}_{i,j}\mathrm{e}^{J_{j}^{*}t}\mathrm{d}t\right\}\mathrm{e}^{-J_{j}^{*}T}.$
There exists a constant $c\geq 0$ such that
$\left\|{\mathrm{e}^{J_{i}t}}\right\|\leq ct^{n_{i}-1}\mathrm{e}^{\operatorname{Re}(\lambda_{i})t},$
$\left\|{\mathrm{e}^{J_{j}^{*}t}}\right\|\leq ct^{n_{j}-1}\mathrm{e}^{\operatorname{Re}(\lambda_{j})t},$
$\left\|{\widetilde{B}_{i,j}}\right\|\leq c$
for all $t\geq 0$. Consequently, we obtain
$\left\|{\int_{0}^{T}\mathrm{e}^{J_{i}t}\widetilde{B}_{i,j}\mathrm{e}^{J_{j}^{*}t}\mathrm{d}t}\right\|$
$\leq\int_{0}^{T}\left\|{\mathrm{e}^{J_{i}t}}\right\|\left\|{\mathrm{e}^{J_{j}^{*}t}}\right\|\left\|{\widetilde{B}_{i,j}}\right\|\mathrm{d}t$
$\leq\int_{0}^{T}c^{3}t^{n_{i}+n_{j}-2}\mathrm{e}^{(\operatorname{Re}(\lambda_{i})+\operatorname{Re}(\lambda_{j}))t}\mathrm{d}t$
$\leq\left\{\begin{aligned} &c^{\prime}\qquad\text{if $\operatorname{Re}(\lambda_{i})+\operatorname{Re}(\lambda_{j})<0$},\\ &c^{\prime}T^{n_{i}+n_{j}-2}\mathrm{e}^{(\operatorname{Re}(\lambda_{i})+\operatorname{Re}(\lambda_{j}))T}\\ &\qquad\text{if $\operatorname{Re}(\lambda_{i})+\operatorname{Re}(\lambda_{j})>0$},\\ &c^{\prime}T^{n_{i}+n_{j}-1}\qquad\text{if $\operatorname{Re}(\lambda_{i})+\operatorname{Re}(\lambda_{j})=0$},\end{aligned}\right$.
for some constant $c^{\prime}\geq 0$. Since
$\left\|{\mathrm{e}^{-J_{j}^{*}T}}\right\|\leq cT^{n_{j}-1}\mathrm{e}^{-\operatorname{Re}(\lambda_{j})T},$
we further obtain
$\left\|{\left\{\int_{0}^{T}\mathrm{e}^{J_{i}t}\widetilde{B}_{i,j}\mathrm{e}^{J_{j}^{*}t}\mathrm{d}t\right\}\mathrm{e}^{-J_{j}^{*}T}}\right\|$
$\leq\left\|{\int_{0}^{T}\mathrm{e}^{J_{i}t}\widetilde{B}_{i,j}\mathrm{e}^{J_{j}^{*}t}\mathrm{d}t}\right\|\left\|{\mathrm{e}^{-J_{j}^{*}T}}\right\|$
[…]
Since $\operatorname{Re}(\lambda_{i})<0<\operatorname{Re}(\lambda_{j})$, the $(i,j)$-block converges to zero in all cases.
Case 4)
Similar to Case 2), the $(i,j)$-block converges to zero.
Case 5)
Since
$\mathrm{e}^{J_{i}t}=\sum_{k=0}^{n_{i}-1}\mathrm{e}^{\lambda_{i}t}\dfrac{t^{k}}{k!}N_{n_{i}}^{k},$
we obtain
$\mathrm{e}^{J_{i}t}\widetilde{B}_{i,j}\mathrm{e}^{J_{j}^{*}t}=\sum_{k=0}^{n_{i}-1}\sum_{\ell=0}^{n_{j}-1}\mathrm{e}^{(\lambda_{i}+\overline{\lambda_{j}})t}\dfrac{t^{k+\ell}}{k!\ell!}N_{n_{i}}^{k}\widetilde{B}_{i,j}(N_{n_{j}}^{*})^{\ell}.$
Consequently, 28 can be written as
$\sum_{k=0}^{n_{i}-1}\sum_{\ell=0}^{n_{j}-1}\left\{\int_{0}^{T}\mathrm{e}^{(\lambda_{i}+\overline{\lambda_{j}})t}\dfrac{t^{k+\ell}}{k!\ell!}\mathrm{d}t\right\}\times\\ D_{i}(T)^{-1}N_{n_{i}}^{k}\widetilde{B}_{i,j}(N_{n_{j}}^{*})^{\ell}D_{j}(T)^{-*}.$ (29)
Since the potentially nonzero entries of $N_{n_{i}}^{k}\widetilde{B}_{i,j}(N_{n_{j}}^{*})^{\ell}$ are confined to its upper-left $(n_{i}-k)\times(n_{j}-\ell)$ block, it follows that
$D_{i}(T)^{-1}N_{n_{i}}^{k}\widetilde{B}_{i,j}(N_{n_{j}}^{*})^{\ell}D_{j}(T)^{-*}$
$=T^{-(k+\ell+1)}\widetilde{D}_{i,k}(T)N_{n_{i}}^{k}\widetilde{B}_{i,j}(N_{n_{j}}^{*})^{\ell}\widetilde{D}_{j,\ell}(T)^{*},$
where
$\widetilde{D}_{i,k}(T)$ $\coloneq\operatorname{diag}(T^{-(n_{i}-k-1)},\ldots,T^{-1},1,0,\ldots,0),$
$\widetilde{D}_{j,\ell}(T)$ $\coloneq\operatorname{diag}(T^{-(n_{j}-\ell-1)},\ldots,T^{-1},1,0,\ldots,0).$
We define $E_{i,k}\coloneq\lim_{T\to\infty}\widetilde{D}_{i,k}(T)$, where $E_{i,k}$ has a $1$ in the $(n_{i}-k,n_{i}-k)$-entry and zeros elsewhere; similarly, we define $E_{j,\ell}\coloneq\lim_{T\to\infty}\widetilde{D}_{j,\ell}(T)$.
Moreover, noting that $\operatorname{Re}(\lambda_{i})=\operatorname{Re}(\lambda_{j})=0$, we obtain
$\left\{\begin{aligned} \left|\int_{0}^{T}\mathrm{e}^{(\lambda_{i}+\overline{\lambda_{j}})t}\dfrac{t^{k+\ell}}{k!\ell!}\mathrm{d}t\right|&\leq cT^{k+\ell}&\text{if $\lambda_{i}\neq\lambda_{j}$},\\ \int_{0}^{T}\mathrm{e}^{(\lambda_{i}+\overline{\lambda_{j}})t}\dfrac{t^{k+\ell}}{k!\ell!}\mathrm{d}t&=\dfrac{T^{k+\ell+1}}{k!\ell!(k+\ell+1)}&\text{if $\lambda_{i}=\lambda_{j}$}\end{aligned}\right$.
for some constant $c\geq 0$. Consequently, the summand of 29 converges, as $T\to\infty$, to
$\left\{\begin{aligned} &0&\text{if $\lambda_{i}\neq\lambda_{j}$},\\ &\dfrac{1}{k!\ell!(k+\ell+1)}E_{i,k}N_{n_{i}}^{k}\widetilde{B}_{i,j}(N_{n_{j}}^{*})^{\ell}E_{j,\ell}^{*}&\text{if $\lambda_{i}=\lambda_{j}$}.\end{aligned}\right$.
Then, $E_{i,k}N_{n_{i}}^{k}\widetilde{B}_{i,j}(N_{n_{k}}^{*})^{\ell}E_{j,\ell}^{*}\in\mathbb{C}^{n_{i}\times n_{j}}$ has a $\widetilde{b}_{i,j}$ in the $(n_{i}-k,n_{j}-\ell)$-entry and zeros elsewhere. Hence, as $T\to\infty$, the $(i,j)$-block converges to $\widetilde{b}_{i,j}C_{i,j}$, where $C_{i,j}$ is defined in 9, when $\lambda_{i}=\lambda_{j}$, whereas it converges to zero when $\lambda_{i}\neq\lambda_{j}$. Furthermore, the central block of $W(B;T)$ of size $n_{0}\times n_{0}$ converges to $W_{0}(B)$.
Case 6)
Noting that $D_{j}(T)=\mathrm{e}^{J_{j}T}$, we can rewrite 28 as
$D_{i}(T)^{-1}\mathrm{e}^{J_{i}T}\left\{\int_{0}^{T}\mathrm{e}^{J_{i}(t-T)}\widetilde{B}_{i,j}\mathrm{e}^{J_{j}^{*}(t-T)}\mathrm{d}t\right\}$
$=D_{i}(T)^{-1}\mathrm{e}^{J_{i}T}\left\{\int_{0}^{T}\mathrm{e}^{-J_{i}t}\widetilde{B}_{i,j}\mathrm{e}^{-J_{j}^{*}t}\mathrm{d}t\right\}.$
Since $\operatorname{Re}(\lambda_{i})=0<\operatorname{Re}(\lambda_{j})$, it follows from Lemma 4 that the integral of the right-hand side converges as $T\to\infty$. Moreover, a direct calculation yields
$D_{i}(T)^{-1}\mathrm{e}^{J_{i}T}$
$=\mathrm{e}^{\lambda_{i}T}\begin{pmatrix}T^{-(n_{i}-\frac{1}{2})}&T^{-(n_{i}-\frac{3}{2})}&\cdots&\dfrac{1}{(n_{i}-1)!}T^{-\frac{1}{2}}\\ &T^{-(n_{i}-\frac{3}{2})}&\cdots&\dfrac{1}{(n_{i}-2)!}T^{-\frac{1}{2}}\\ &&\ddots&\vdots\\ &&&T^{-\frac{1}{2}}\end{pmatrix}$,
and thus $D_{i}(T)\mathrm{e}^{J_{i}T}\to 0$ as $T\to\infty$. Hence, the $(i,j)$-block converges to zero.
Case 7)
Similar to Case 3), the $(i,j)$-block converges to zero.
Case 8)
Similar to Case 6), the $(i,j)$-block converges to zero.
Case 9)
Noting that $D_{i}(T)=\mathrm{e}^{J_{i}T}$ and $D_{j}(T)=\mathrm{e}^{J_{j}T}$, 28 reduces to
$\int_{0}^{T}\mathrm{e}^{J_{i}(t-T)}\widetilde{B}_{i,j}\mathrm{e}^{J_{j}^{*}(t-T)}\mathrm{d}t$
$=\int_{0}^{T}\mathrm{e}^{-J_{i}t}\widetilde{B}_{i,j}\mathrm{e}^{-J_{j}^{*}t}\mathrm{d}t.$
Since $\operatorname{Re}(\lambda_{i}),\operatorname{Re}(\lambda_{j})>0$, it follows from Lemma 4 that the integral of the right-hand side converges, as $T\to\infty$, to the unique solution $W_{+}^{(i,j)}(B)\in\mathbb{C}^{n_{i}\times n_{j}}$ satisfying
$(-J_{i})W_{+}^{(i,j)}(B)+W_{+}^{(i,j)}(B)(-J_{j}^{*})+\widetilde{B}_{i,j}=0.$
Then, the lower-right block of $W(B;T)$ of size $n_{+}\times n_{+}$ converges to
$\begin{pmatrix}W_{+}^{(\gamma+1,\gamma+1)}(B)&\cdots&W_{+}^{(\gamma+1,m)}(B)\\ \vdots&\ddots&\vdots\\ W_{+}^{(m,\gamma+1)}(B)&\cdots&W_{+}^{(m,m)}(B)\end{pmatrix}$,
where we set $\gamma\coloneq m_{-}+m_{0}$ to simplify the indexing. The matrix coincides with $W_{+}(B)$.
Summarizing the discussion so far, $W(B;T)$ converges to $W_{\infty}(B)$, completing the proof. $\Box$
Appendix B Proof of Lemmas 2 and 3
First, we prove Lemma 2.
Proof (Lemma 2)
Let $y\in\mathbb{R}^{n}$ be a nonzero vector. Since $\mathrm{e}^{A^{\top}t}$ is nonsingular for all $t$, $\mathrm{e}^{A^{\top}t}y\neq 0$. Therefore, since $\operatorname{diag}(p)$ is positive definite,
$y^{\top}\widetilde{W}(p;T)y$ $=\int_{0}^{T}y^{\top}\mathrm{e}^{At}\operatorname{diag}(p)\mathrm{e}^{A^{\top}t}y\mathrm{d}t$
$>0.$
This proves that $\widetilde{W}(p;T)$ is positive definite. $\Box$
Next, to prove Lemma 3, we employ the following lemma.
Lemma 5 ([14, Theorem 12])
Let $C\in\mathbb{C}^{\ell\times\ell}$ be Hurwitz and $P\in\mathbb{C}^{\ell\times\ell}$ be Hermitian positive definite. Then, the Lyapunov equation
$CY+YC^{*}+P=0$
admits a unique solution $Y\in\mathbb{C}^{\ell\times\ell}$, which is Hermitian positive definite.
Proof (Lemma 3)
From 12, it suffices to prove that $W_{-}(p),W_{0}(p)$, and $W_{+}(p)$ are positive definite. Let $B\coloneq\operatorname{diag}(p),\widetilde{B}\coloneq P^{-1}BP^{-*}$, and partition $\widetilde{B}$ in the same way as 6. Then, the matrices $\widetilde{B}_{-,-},\widetilde{B}_{0,0}$, and $\widetilde{B}_{+,+}$ are positive definite since they are principal submatrices of $\widetilde{B}$, which is positive definite.
From 7, the matrix $W_{-}(p)$ is the unique solution to
$J_{-}W_{-}(p)+W_{-}(p)J_{-}^{*}+\widetilde{B}_{-,-}=0,$
and it follows from Lemma 5 that $W_{-}(p)$ is positive definite. Similarly, we can prove that $W_{+}(p)$ is positive definite.
We next prove that $W_{0}(p)$ is positive definite. Since $W_{0}(p)$ is a block-diagonal matrix from 11, it suffices to show that $W_{0}^{(\ell)}(p)$ is positive definite for $\ell=1,\ldots,m_{0}^{\prime}$. To simplify the indexing, let $\alpha=m_{-}+m_{0}^{(1)}+\cdots+m_{0}^{(\ell-1)},\ \beta=m_{-}+m_{0}^{(1)}+\cdots+m_{0}^{(\ell)}$. Let $e^{(i)}\coloneq(0,\ldots,0,1)^{\top}\in\mathbb{R}^{n_{i}\times n_{i}}$. Then, a direct calculation yields
$C_{i,j}=\int_{0}^{1}\mathrm{e}^{N_{i}t}e^{(i)}e^{(j)}{}^{\top}\mathrm{e}^{N_{j}^{\top}t}\mathrm{d}t.$
Furthermore, let
$\widetilde{B}_{0,0}^{(\ell)}\coloneq\begin{pmatrix}\widetilde{b}_{\alpha+1,\alpha+1}&\cdots&\widetilde{b}_{\alpha+1,\beta}\\ \vdots&\ddots&\vdots\\ \widetilde{b}_{\beta,\alpha+1}&\cdots&\widetilde{b}_{\beta,\beta}\end{pmatrix}\in\mathbb{C}^{m_{0}^{(\ell)}\times m_{0}^{(\ell)}},$
which is positive definite since it is a principal submatrix of $\widetilde{B}_{0,0}$,
$E^{(\ell)}$ $\coloneq\begin{pmatrix}e^{(\alpha+1)}&0&\cdots&0&0\\ 0&e^{(\alpha+2)}&\cdots&0&0\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&\cdots&e^{(\beta-1)}&0\\ 0&0&\cdots&0&e^{(\beta)}\end{pmatrix}$
$\in\mathbb{R}^{n_{0}^{(\ell)}\times m_{0}^{(\ell)}},$
and
$J_{0}^{(\ell)}\coloneq\begin{pmatrix}J_{\alpha+1}\\ &\ddots\\ &&J_{\beta}\end{pmatrix}\in\mathbb{C}^{n_{0}^{(\ell)}\times n_{0}^{(\ell)}}.$
By further calculation, it follows from 10 that
$W_{0}^{(\ell)}(p)=\int_{0}^{1}\mathrm{e}^{J_{0}^{(\ell)}t}E^{(\ell)}\widetilde{B}_{0,0}^{(\ell)}E^{(\ell)}{}^{\top}\mathrm{e}^{J_{0}^{(\ell)}{}^{*}t}\mathrm{d}t.$
Let $y\in\mathbb{C}^{n_{0}^{(\ell)}\times n_{0}^{(\ell)}}$ be a nonzero vector, and partition it into blocks as $y=(y^{(\alpha+1)}{}^{\top},\ldots,y^{(\beta)}{}^{\top})^{\top}$. Then, let us consider
$y^{*}W_{0}^{(\ell)}(p)y=\int_{0}^{1}y^{*}\mathrm{e}^{J_{0}^{(\ell)}t}E^{(\ell)}\widetilde{B}_{0,0}^{(\ell)}E^{(\ell)}{}^{\top}\mathrm{e}^{J_{0}^{(\ell)}{}^{*}t}y\mathrm{d}t.$
Since $y$ is nonzero,
$y^{*}\mathrm{e}^{J_{0}^{(\ell)}t}E^{(\ell)}=(y^{(\alpha+1)}{}^{*}\mathrm{e}^{J_{\alpha+1}t}e^{(\alpha+1)},\cdots,y^{(\beta)}{}^{*}\mathrm{e}^{J_{\beta}t}e^{(\beta)})$
is also a nonzero vector for all $t\in[0,1]$. Since $\widetilde{B}_{0,0}^{(\ell)}$ is positive definite,
$y^{*}\mathrm{e}^{J_{0}^{(\ell)}t}E^{(\ell)}\widetilde{B}_{0,0}^{(\ell)}E^{(\ell)}{}^{\top}\mathrm{e}^{J_{0}^{(\ell)}{}^{*}t}y>0$
holds for all $t\in[0,1]$. This proves that $W_{0}^{(\ell)}(p)$ is positive definite, completing the proof. $\Box$
Appendix C Proof of Theorems 2 and 3
[…]
Let $\{f^{(k)}\}_{k\in\mathbb{N}}$ be a sequence of functions $f^{(k)}\colon\mathbb{R}^{n-1}\to\mathbb{R}\cup\{\infty\}$. We say that $\{f^{(k)}\}_{k\in\mathbb{N}}$ is equi level-bounded if there exists a bounded set $B\subset\mathbb{R}^{n-1}$ such that $\mathrm{lev}_{\leq\alpha}f^{(k)}\subset B$ for all $k$. We say that $\{f^{(k)}\}_{k\in\mathbb{N}}$ epi-converges to $f$ if, for every $p^{\prime}\in\mathbb{R}^{n-1}$,
[…]
We employ the following three propositions.
Proposition 1 ([21, Theorem 1.9])
Assume that $f\colon\mathbb{R}^{n-1}\to\mathbb{R}\cup\{\infty\}$ is proper, lower semicontinuous, and level-bounded. Then, $\operatorname*{arg\,min}f$ is nonempty and compact, and the minimum value of $f$ is finite.
Proposition 2 ([21, Theorem 7.17])
[…]
Proposition 3 ([21, Theorem 7.33])
Assume that $\{f^{(k)}\}_{k\in\mathbb{N}}$ is an equi level-bounded sequence of functions $f^{(k)}\colon\mathbb{R}^{n-1}\to\mathbb{R}\cup\{\infty\}$, that $\{f^{(k)}\}_{k\in\mathbb{N}}$ epi-converges to $f$, and that $f^{(k)}$ and $f$ are lower semicontinuous and proper. Then, the sequence $\{\widetilde{p}^{(k)}\}_{k\in\mathbb{N}}$ such that $p^{(k)}\in\operatorname*{arg\,min}f^{(k)}$ is bounded, and all its cluster points lie in $\operatorname*{arg\,min}f$. In particular, when $\operatorname*{arg\,min}f=\{\widetilde{p}^{*}\}$, the sequence $\{\widetilde{p}^{(k)}\}_{k\in\mathbb{N}}$ converges to $\widetilde{p}^{*}$.
Furthermore, we prepare the following lemma to prove Theorem 3.
Lemma 6
Let $C_{1},C_{2}\in\mathbb{C}^{\ell\times\ell}$ be Hermitian positive definite matrices, and $\mu>0$ is the minimum eigenvalue of $C_{2}$. Then,
$\operatorname{tr}(C_{1}C_{2})\geq\mu\operatorname{tr}(C_{1})$
holds.
Proof
Since $C_{2}\succeq\mu I_{\ell}$, we obtain $C_{1}^{1/2}C_{2}C_{1}^{1/2}\succeq\mu C_{1}$. Therefore,
$\operatorname{tr}(C_{1}C_{2})=\operatorname{tr}(C_{1}^{1/2}C_{2}C_{1}^{1/2})\geq\mu\operatorname{tr}(C_{1})$
holds. $\Box$
Then, we can prove Theorems 2 and 3.
Proof (Theorem 2)
Let $\{T_{k}\}_{k\in\mathbb{N}}$ be any sequence such that $T_{k}>0$ and $T_{k}\to\infty$ as $k\to\infty$. It suffices to prove that $p_{\mathrm{VCS}}^{T_{k}}\to p_{\mathrm{VCS}}^{\infty}$ as $k\to\infty$.
Define $g^{(k)}\colon\mathbb{R}^{n-1}\to\mathbb{R}\cup\{\infty\}$ by
$g^{(k)}(\widetilde{p})\coloneq\left\{\begin{aligned} &-\log\det W(\varphi(\widetilde{p});T_{k})\\ &\qquad\text{if $\widetilde{p}\in\widetilde{\Delta},\ W(\varphi(\widetilde{p});T_{k})\succ O$,}\\ &\infty\quad\text{otherwise},\end{aligned}\right$.
and define $g\colon\mathbb{R}^{n-1}\to\mathbb{R}\cup\{\infty\}$ by
[…]
Then, $\{g^{(k)}\}_{k\in\mathbb{N}}$ is a sequence of convex functions and equi level-bounded since $\mathrm{lev}_{\leq\alpha}g^{(k)}\subset\widetilde{\Delta}$ for every $\alpha\in\mathbb{R}$. Similarly, $g$ is convex and level-bounded. If follows from Lemma 2 that $g^{(k)}$ is proper. Furthermore, it follows from Lemma 3 that $\operatorname*{int}\operatorname*{dom}g=\operatorname*{int}\widetilde{\Delta}\neq\emptyset$.
Next, we prove that $g$ is lower semicontinuous, that is, $\mathrm{lev}_{\leq\alpha}g$ is closed for every $\alpha\in\mathbb{R}$. Let $\widetilde{p}^{(k)}\in\mathrm{lev}_{\leq\alpha}g$ with $\widetilde{p}^{(k)}\to\widetilde{p}\in\mathbb{R}^{n-1}$ as $k\to\infty$. Since $\det W_{\infty}(\varphi(\widetilde{p}^{(k)}))\geq\mathrm{e}^{-\alpha}$, by continuity, we obtain $\det W_{\infty}(\varphi(\widetilde{p}))\geq\mathrm{e}^{-\alpha}$ and $W_{\infty}(\varphi(\widetilde{p}))\succ O$. Since $\widetilde{\Delta}$ is closed, we obtain $\widetilde{p}\in\Delta$; hence $g(\widetilde{p})\leq\alpha$, implying $\widetilde{p}\in\mathrm{lev}_{\leq\alpha}g$. This shows that $\mathrm{lev}_{\leq\alpha}g$ is closed. Similarly, we can prove that $g^{(k)}$ is lower semicontinuous.
[…]
Proof (Theorem 3)
Let $\{T_{k}\}_{k\in\mathbb{N}}$ be any sequence such that $T_{k}>0$ and $T_{k}\to\infty$ as $k\to\infty$. It suffices to prove that $p_{\mathrm{AECS}}^{T_{k}}\to p_{\mathrm{AECS}}^{\infty}$ as $k\to\infty$.
Define $h^{(k)}\colon\mathbb{R}^{n-1}\to\mathbb{R}\cup\{\infty\}$ by
$h^{(k)}(\widetilde{p})\coloneq\left\{\begin{aligned} &\operatorname{tr}\left(R^{-*}D(T_{k})^{-*}W(\varphi(\widetilde{p});T_{k})^{-1}D(T_{k})^{-1}R^{-1}\right)\\ &\qquad\text{if $\widetilde{p}\in\widetilde{\Delta},\ W(\varphi(\widetilde{p});T_{k})\succ O$,}\\ &\infty\quad\text{otherwise},\end{aligned}\right$.
and define $h\colon\mathbb{R}^{n-1}\to\mathbb{R}\cup\{\infty\}$ by
[…]
Then, $\{h^{(k)}\}_{k\in\mathbb{N}}$ is a sequence of convex functions and equi level-bounded since $\mathrm{lev}_{\leq\alpha}h^{(k)}\subset\widetilde{\Delta}$ for every $\alpha\in\mathbb{R}$. Similarly, $h$ is convex and level-bounded. If follows from Lemma 2 that $h^{(k)}$ is proper. Furthermore, it follows from Lemma 3 that $\operatorname*{int}\operatorname*{dom}h=\operatorname*{int}\widetilde{\Delta}\neq\emptyset$.
Next, we prove that $h$ is lower semicontinuous. Let $(R^{-1}R^{-*})_{-}$ be the upper-left block of $R^{-1}R^{-*}$ of size $n_{-}\times n_{-}$. Then, we obtain
$\operatorname{tr}\left(R^{-*}\begin{pmatrix}W_{-}(\varphi(\widetilde{p}))^{-1}\\ &O\end{pmatrix}R^{-1}\right)$
$=\operatorname{tr}\left(W_{-}(\varphi(\widetilde{p}))^{-1}(R^{-1}R^{-*})_{-}\right).$
Furthermore, since $R^{-1}R^{-*}$ is Hermitian positive definite, $(R^{-1}R^{-*})_{-}$ is also Hermitian positive definite, and let $\mu>0$ denote its minimum eigenvalue.
Let $\alpha\in\mathbb{R}$ be arbitray, and let $\widetilde{p}^{(k)}\in\mathrm{lev}_{\leq\alpha}h$ with $\widetilde{p}^{(k)}\to\widetilde{p}\in\mathbb{R}^{n-1}$ as $k\to\infty$. Since $\widetilde{\Delta}$ is closed, we obtain $\widetilde{p}\in\Delta$. Suppose that $W_{-}(\varphi(\widetilde{p}))\not\succ O$. Then, $\operatorname{tr}\left(W_{-}(\varphi(\widetilde{p}^{(k)}))^{-1}\right)\to\infty$. It follows from Lemma 6 that
$h(\widetilde{p}^{(k)})$ $=\operatorname{tr}\left(W_{-}(\varphi(\widetilde{p}^{(k)}))^{-1}(R^{-1}R^{-*})_{-}\right)$
$\geq\mu\operatorname{tr}\left(W_{-}(\varphi(\widetilde{p}^{(k)}))^{-1}\right)\to\infty.$
However, this contradicts that $h(\widetilde{p}^{(k)})\leq\alpha$ for all $k$. Hence, $W_{-}(\varphi(\widetilde{p}))\succ O$.
By continuity, we obtain
$h(\widetilde{p})=\operatorname{tr}\left(W_{-}(\varphi(\widetilde{p}))^{-1}(R^{-1}R^{-*})_{-}\right)\leq\alpha,$
and therefore $\widetilde{p}\in\mathrm{lev}_{\leq\alpha}h$. This shows that $\mathrm{lev}_{\leq\alpha}h$ is closed, implying that $h$ is lower semicontinuous. Similarly, we can prove that $h^{(k)}$ is lower semicontinuous. The proof proceeds in the same manner as that of Theorem 2. $\Box$
Appendix D Proof of Theorems 4 and 5
To prove Theorems 4 and 5, we prepare the following lemmas, which can be proved in the same manner as [23, Theorems 1 and 3], although Problems 2 and 6 have slightly different forms.
Lemma 7
Assume that $W_{\infty,1},\ldots,W_{\infty,n}$ are linearly independent over $\mathbb{R}$. Then, Problem 5 admits a unique optimal solution.
Lemma 8
Assume that $W_{-,1},\ldots,W_{-,n}$ are linearly independent over $\mathbb{R}$. Then, Problem 6 admits a unique optimal solution.
Proof (Theorem 4)
Let $x=(x_{1},\ldots,x_{n})^{\top}\in\mathbb{R}^{n}$ be any vector such that $W_{\infty}(x)=0$. Then, from Lemma 7, it suffices to prove $x=0$.
Let $\widetilde{X}\coloneq Q^{-1}XQ^{-*}$ and partition $\widetilde{X}$ into blocks in following two ways:
$\widetilde{X}=\begin{pmatrix}\widetilde{X}_{1,1}&\cdots&\widetilde{X}_{1,m}\\ \vdots&\ddots&\vdots\\ \widetilde{X}_{m,1}&\cdots&\widetilde{X}_{m,m}\end{pmatrix}\\ =\begin{pmatrix}\widetilde{X}_{-,-}&\widetilde{X}_{-,0}&\widetilde{X}_{-,+}\\ \widetilde{X}_{0,-}&\widetilde{X}_{0,0}&\widetilde{X}_{0,+}\\ \widetilde{X}_{+,-}&\widetilde{X}_{+,0}&\widetilde{X}_{+,+}\end{pmatrix}$. (30)
Furthermore, we denote the $(n_{i},n_{j})$-entry of $X_{i,j}$ by $\widetilde{x}_{i,j}$, which is the lower-right entry, and for $\ell=1,\ldots,m_{0}^{\prime}$, let
$\widetilde{X}_{0,0}^{(\ell)}\coloneq\begin{pmatrix}\widetilde{x}_{\alpha+1,\alpha+1}&\cdots&\widetilde{x}_{\alpha+1,\beta}\\ \vdots&\ddots&\vdots\\ \widetilde{x}_{\beta,\alpha+1}&\cdots&\widetilde{x}_{\beta,\beta}\end{pmatrix}\in\mathbb{C}^{m_{0}^{(\ell)}\times m_{0}^{(\ell)}},$
where we set $\alpha=m_{-}+m_{0}^{(1)}+\cdots+m_{0}^{(\ell-1)},\ \beta=m_{-}+m_{0}^{(1)}+\cdots+m_{0}^{(\ell)}$ to simplify the indexing.
Then, under the assumption that $W_{\infty}(x)=O$, it follows that $\widetilde{X}_{-,-}=O,\widetilde{X}_{+,+}=O$, and $\widetilde{X}_{0,0}^{(\ell)}=O\ (\ell=1,\ldots,m_{0}^{\prime})$. Otherwise, the argument in Section III-B would yield $W_{\infty}(x)\neq 0$. Furthermore, we obtain
$\widetilde{X}_{-,-}=\sum_{i=1}^{n}x_{i}q_{-,i}q_{-,i}^{*},$
$\widetilde{X}_{+,+}=\sum_{i=1}^{n}x_{i}q_{+,i}q_{+,i}^{*},$
$\widetilde{X}_{0,0}^{(\ell)}=\sum_{i=1}^{n}x_{i}q_{0,i}^{(\ell)}q_{0,i}^{(\ell)}{}^{*}\quad(\ell=1,\ldots,m_{0}^{\prime}).$
By vectorizing the expression, we obtain $M_{\mathrm{VCS}}x=0$. Since $\operatorname{rank}M_{\mathrm{VCS}}=n$, it follows that $x=0$, which completes the proof. $\Box$
Proof (Theorem 5)
Let $x=(x_{1},\ldots,x_{n})^{\top}\in\mathbb{R}^{n}$ be any vector such that $W_{-}(x)=0$. Then, from Lemma 8, it suffices to prove $x=0$. Let $\widetilde{X}\coloneq Q^{-1}XQ^{-*}$ and partition $\widetilde{X}$ in the same way as 30. By an argument analogous to that of Theorem 4, we obtain $\widetilde{X}_{-,-}=O$. Proceeding as in the proof of Theorem 4, the assumption then implies that $x=0$, which completes the proof. $\Box$
Appendix E Uniqueness of finite-horizon controllability scores
In [23], an example is presented in which the optimal solutions to Problems 1 and 2 are not unique for some $T>0$. As noted in Section IV-C, the uniqueness of the optimal solutions is essential when interpreting controllability scores as network centrality. In this section, we summarize existing results on the uniqueness of the controllability score on a finite time horizon. We further present new results that encompass the existing ones and are applied to the Laplacian dynamics.
For Hurwitz matrices, it has been shown that the optimal solution is unique for all $T>0$.
Proposition 4 ([23, Theorem 4])
Assume that $A$ is Hurwitz. Then, Problems 1 and 2 each admit a unique optimal solution for all $T>0$.
An important class of examples that does not satisfy this assumption is given by graph Laplacians. Since a graph Laplacian has zero as an eigenvalue, it fails to satisfy the assumption. For graph Laplacians of undirected connected graphs, the following result has been established.
Proposition 5 ([23, Theorem 5])
Let $L$ be the Laplacian matrix of an undirected connected graph, and let A=LA=-L. Then, Problems 1 and 2 each admit a unique optimal solution for all $T>0$.
Without restricting the class of systems, the following result has been established for arbitrary matrices.
Proposition 6 ([22, Theorem 1])
Let $A$ be an arbitrary matrix. Then, Problems 1 and 2 each admit a unique optimal solution for almost all $T>0$.
We present the following result, which encompasses all of the above results.
Theorem 7
Let
$\Theta\coloneq\left\{\theta\in\mathbb{R}\setminus\{0\}\,\middle|\,\text{$\lambda_{i}+\lambda_{j}=\theta\sqrt{-1}$ for some $i,j$}\right\},$
and
$\Theta^{\prime}\coloneq\left\{\left|\dfrac{2\ell\pi}{\theta}\right|\in\mathbb{R}_{>0}\,\middle|\,\theta\in\Theta,\ell\in\mathbb{Z}\setminus\{0\}\right\}.$
Then, Problems 1 and 2 each admit a unique optimal solution for $T\not\in\Theta^{\prime}$.
To prove Theorem 7, we introduce some notation and prepare several lemmas. For $Y=(y_{1},\ldots,y_{\ell^{\prime}}),\ y_{i}\in\mathbb{R}^{\ell}$, we define
$\operatorname{vec}(Y)\coloneq(y_{1}^{\top},\ldots,y_{\ell^{\prime}}^{\top})^{\top}\in\mathbb{R}^{\ell\ell^{\prime}}.$
Then, it is known that the following equality holds [10]:
$\operatorname{vec}(CYC^{\prime\top})=(C^{\prime}\otimes C)\operatorname{vec}(Y).$ (31)
Let $K\coloneq I_{n}\otimes A+A\otimes I_{n}\in\mathbb{R}^{n^{2}\times n^{2}}$. The following result is known for the spectrum of $K$ (see, e.g., [10, Appendix B.14]).
Lemma 9
Every eigenvalue of $K$ is of the form $\lambda_{i}+\lambda_{j}$ for some $i,j$.
We employ the following spectral mapping theorem.
Lemma 10 ([10, Theorem 1.13])
Let $f\colon\mathbb{C}\to\mathbb{C}$ be a holomorphic function. Then, every eigenvalue of $f(K)$ is of the form $f(\mu)$, where $\mu$ is an eigenvalue of $K$.
Moreover, we employ the following lemma, which can be proved in the same way as in [23, Theorems 1 and 3].
Lemma 11
Let $T>0$ be arbitrary. If $\widetilde{W}(x;T)=0$ for $x\in\mathbb{R}^{n}$ implies $x=0$, then, Problems 1 and 2 each admit a unique optimal solution for that $T$.
Proof (Theorem 7)
Let us consider a linear map
$F_{T}\colon\mathbb{R}^{n\times n}\to\mathbb{R}^{n\times n},\quad B\mapsto\widetilde{W}(B;T).$ (32)
Assume that $F_{T}(B)=0$ implies $B=0$. In particular, if $\widetilde{W}(x;T)=0$, then $x=0$, and by Lemma 11, the optimal solution is unique for that $T$. We therefore investigate whether $F_{T}$ has zero as an eigenvalue.
It follows from 31 that
$\operatorname{vec}\left(\widetilde{W}(B;T)\right)$ $=\int_{0}^{T}\operatorname{vec}\left(\mathrm{e}^{At}B\mathrm{e}^{A^{\top}t}\right)\mathrm{d}t$
$=\left(\int_{0}^{T}\mathrm{e}^{At}\otimes\mathrm{e}^{At}\mathrm{d}t\right)\operatorname{vec}(B)$
$=\left(\int_{0}^{T}\mathrm{e}^{Kt}\mathrm{d}t\right)\operatorname{vec}(B).$
Thus, by vectorization, the map $F_{T}$ can be identified with the linear map
$\mathbb{R}^{n^{2}}\to\mathbb{R}^{n^{2}},\quad\operatorname{vec}(B)\mapsto\left(\int_{0}^{T}\mathrm{e}^{Kt}\mathrm{d}t\right)\operatorname{vec}(B).$ (33)
Let
[…]
which is a holomorphic function. By identifying 32 and 33, we have $F_{T}=f_{T}(K)$. Consequently, from Lemmas 9 and 10, every eigenvalue of $F_{T}$ is of the form $f_{T}(\lambda_{i}+\lambda_{j})$ for some $i,j$.
Assume $f_{T}(\lambda_{i}+\lambda_{j})=0$ for some $i,j$. Then, $\lambda_{i}+\lambda_{j}=\dfrac{2\ell\pi\sqrt{-1}}{T}$ for some $\ell\in\mathbb{Z}\setminus\{0\}$. In this case, $\lambda_{i}+\lambda_{j}$ is purely imaginary, and we can write $\lambda_{i}+\lambda_{j}=\theta\sqrt{-1}$ for some $\theta\neq 0$. Consequently, $T=\dfrac{2\ell\pi}{\theta}=\left|\dfrac{2\ell\pi}{\theta}\right|$, and $T\in\Theta^{\prime}$.
Therefore, if $T\not\in\Theta^{\prime}$, then $F_{T}$ does not have zero as an eigenvalue. It follows from Lemma 11 that Problems 1 and 2 each admit a unique optimal solution for this $T$. This completes the proof. $\Box$
It follows that values of $T$ for which the optimal solution is not unique are confined to $\Theta^{\prime}$. Since $\Theta^{\prime}$ is at most countable, the optimal solution is unique for almost all $T>0$. In other words, Theorem 7 can be understood as characterizing the exceptional values of $T>0$ for which the optimal solution fails to be unique in the result established by Proposition 6.
Moreover, although Proposition 5 is stated for Laplacians of undirected connected graphs, the same uniqueness property is shown to hold for directed graphs as well.
Corollary 1
Let $L$ be the Laplacian matrix of a directed graph, and let A=LA=-L. Then, Problems 1 and 2 each admit a unique optimal solution for all $T>0$.
Proof
It follows from Example 2 that $\Theta=\emptyset$. Therefore, $\Theta^{\prime}=\emptyset$ also holds, and this completes the proof. $\Box$
Acknowledgment
This work was supported by JST PRESTO, Japan, Grant Number JPMJPR25K4.

### v2

@@
-Controllability score for linear time-invariant systems on an infinite time horizon
+Infinite-horizon controllability scores for linear time-invariant systems
 Abstract
-We introduce a scaled controllability Gramian that can be computed reliably even for unstable systems.
-Using this scaled Gramian, we reformulate the controllability scoring problems into equivalent but numerically stable optimization problems.
-Their optimal solutions define dynamics-aware network centrality measures, referred to as the volumetric controllability score (VCS) and the average energy controllability score (AECS).
-We then formulate controllability scoring problems on an infinite time horizon.
-Under suitable assumptions, we prove that the resulting VCS and AECS are unique and that the finite-horizon scores converge to them.
+We introduce a numerically stable reformulation of controllability scoring based on a scaled controllability Gramian, which remains reliably computable even for unstable systems.
+The resulting optimization problems define dynamics-aware network centrality measures, referred to as the volumetric controllability score (VCS) and the average energy controllability score (AECS).
+Building on this stable reformulation, we derive the corresponding infinite-horizon problems, develop an algorithm to solve them, and highlight computational advantages over their finite-horizon counterparts.
+Under suitable assumptions, we prove that the infinite-horizon VCS and AECS are unique and that the finite-horizon scores converge to them.
 We further show that VCS and AECS can differ markedly in this limit, because VCS enforces controllability of the full system, whereas AECS accounts only for the stable modes.
-Finally, using Laplacian dynamics as a representative example, we present numerical experiments that illustrate this convergence.
+Finally, numerical experiments on Laplacian dynamics illustrate this convergence.
 Index Terms:
@@
 Under mild conditions, controllability scores are defined as the unique optimal solutions of convex optimization problems, thereby assigning each node a quantitative measure of its importance for controlling the network dynamics.
-The controllability scores are designed to address limitations of existing controllability-based frameworks.
-Prior approaches can be broadly categorized into structural controllability [15] and quantitative controllability measures [18].
-Structural controllability depends only on topology and enables efficient graph-theoretic analysis without identifying edge weights [16, 17], but it does not preclude prohibitively large control-energy requirements [30].
-Quantitative measures incorporate edge weights and time horizons to capture energy-related aspects of controllability; building on this viewpoint, controllability scores evaluate node importance while accounting for control effort.
-Existing Gramian-based centralities such as volumetric controllability centrality (VCE) and average control energy centrality (ACE) may still be misaligned with quantitative controllability because they rely on the rank or pseudoinverse of a single-input controllability Gramian [23].
-By contrast, controllability scores are defined by introducing virtual actuation at all state nodes and optimizing the input allocation, rather than relying on a single-input controllability Gramian.
+The controllability scores address limitations of existing controllability-based frameworks, which can be broadly categorized into structural controllability [16] and quantitative controllability measures [19].
+Structural controllability depends only on network topology and enables graph-theoretic analysis without identifying edge weights [17, 18], but it does not preclude prohibitively large control-energy requirements [31].
+Quantitative measures incorporate edge weights and time horizons to capture energy-related aspects of controllability; in this spirit, controllability scores quantify node importance while accounting for control effort.
+Existing Gramian-based centralities such as VCE and ACE introduced in [26] may still be misaligned with quantitative controllability because they rely on the rank or pseudoinverse of a single-input controllability Gramian [24].
+By contrast, controllability scores are defined by introducing virtual actuation at all state nodes and optimizing the input allocation.
 Subsequent studies established uniqueness properties [24, 23], demonstrated applicability to human brain networks [23], and incorporated constraints on actuated nodes [25].
-In Section II-A, we define two controllability scores (VCS and AECS) as optimal solutions to distinct controllability-Gramian–based optimization problems.
-This finite-horizon formulation requires specifying a terminal time $T>0$.
-However, the choice of $T$ has not been systematically addressed in the existing literature.
-It has been reported in [22] that when $T$ is small, the resulting scores are dominated by short-term dynamics and fail to capture meaningful system-wide properties, whereas for sufficiently large $T$, they reflect the network-structured dynamics of the system.
-These observations suggest taking $T$ large.
-A natural extreme is $T=\infty$, which eliminates dependence on the system time scale and on the researcher’s subjective choice, and is therefore desirable from the viewpoint of reproducibility.
-However, except for the case where $A$ is Hurwitz (i.e., all its eigenvalues have strictly negative real parts), the controllability Gramian diverges as $T\to\infty$, so the scores are no longer well defined.
-Moreover, even for large but finite $T$, the Gramian can become ill-conditioned, making objective-function evaluation numerically unstable and potentially preventing accurate computation.
-This paper provides a solution to this issue.
+In Section II-A, we define two controllability scores, VCS and AECS, as optimal solutions to distinct controllability-Gramian–based optimization problems on a finite horizon.
+This formulation requires choosing a terminal time $T>0$.
+While existing work has not systematically addressed this choice, [23] reports that small $T$ yields scores dominated by short-term dynamics, whereas sufficiently large $T$ reveals network-structured, system-wide behavior, suggesting that $T$ should be large.
+A natural extreme is $T=\infty$, which removes the need to tune $T$ and improves reproducibility.
+However, unless $A$ is Hurwitz (i.e., all its eigenvalues have strictly negative real parts), the controllability Gramian diverges as $T\to\infty$, making the scores ill-defined.
+Moreover, even for large but finite $T$, the Gramian can be ill-conditioned, leading to numerically unstable evaluation of the objective and inaccurate computation.
+This paper addresses these issues.
 The contributions of this paper are summarized as follows.
@@
   2. 2.
-    By taking the limit $T\to\infty$ in the reformulated problems, we obtain controllability scores on an infinite time horizon, and prove, under suitable assumptions, that the finite-horizon scores converge to them.
+    By taking the limit $T\to\infty$ in the reformulated problems, we obtain controllability scores on an infinite time horizon, and prove, under suitable assumptions, that these scores are unique and that the finite-horizon scores converge to them.
 This provides a principled choice of the terminal time, namely $T=\infty$, and also offers an explanation for the empirical observation reported in [23] that, although the controllability Gramian diverges for unstable systems, the controllability scores nevertheless converge to a certain value as $T\to\infty$.
@@
   4. 4.
-    Under suitable assumptions, we show that the scaled controllability Gramian can be computed in a numerically stable manner using a real Schur decomposition and the solution of a Sylvester equation.
-This enables the stable numerical evaluation of the infinite-horizon controllability scores.
+    Under suitable assumptions, the infinite-horizon scaled controllability Gramian—and hence the infinite-horizon controllability scores—admits a numerically stable computation using the real Schur decomposition, followed by solving Sylvester and Lyapunov equations.
+Unlike the finite-horizon case, which typically requires numerical integration with a computational cost growing with $T$, the infinite-horizon formulation relies on solving Lyapunov equations with a cost independent of $T$.
+Moreover, this computation can be further accelerated using efficient approximation algorithms such as CF-ADI [15], making the infinite-horizon approach computationally advantageous.
 The rest of this paper is organized as follows.
@@
 The sets of all natural numbers, all real numbers, and all complex numbers are denoted by $\mathbb{N}$, $\mathbb{R}$, and $\mathbb{C}$, respectively.
+The symbol $\sqrt{-1}$ denotes the imaginary unit.
 Let $e_{i}\coloneq(0,\ldots,0,1,0,\ldots,0)^{\top}\in\mathbb{R}^{n}$ be a standard vector whose $i$-th entry is $1$ and other entries are 0.
@@
 Let $O$ denote a zero matrix of a suitable size.
-Let $\Delta$ and $\widetilde{\Delta}$ be the standard simplices in $\mathbb{R}^{n}$ and $\mathbb{R}^{n-1}$, respectively, i.e.,
-$\Delta\coloneq\left\{(p_{i})\in\mathbb{R}^{n}\,\middle|\,\begin{gathered}p_{i}\geq 0\ (i=1,\ldots,n),\\ \textstyle\sum_{i=1}^{n}p_{i}=1\end{gathered}\right\},$
-$\widetilde{\Delta}\coloneq\left\{(\widetilde{p}_{i})\in\mathbb{R}^{n-1}\,\middle|\,\begin{gathered}\widetilde{p}_{i}\geq 0\ (i=1,\ldots,n-1),\\ \textstyle\sum_{i=1}^{n-1}\widetilde{p}_{i}\leq 1\end{gathered}\right\}.$
-The mapping $\varphi\colon\widetilde{\Delta}\to\Delta$ defined by
-$\varphi(\widetilde{p}_{1},\ldots,\widetilde{p}_{n-1})=(\widetilde{p}_{1},\ldots,\widetilde{p}_{n-1},1-\sum_{i=1}^{n-1}\widetilde{p}_{i})^{\top}$
-is a homeomorphism; its inverse is given by
-$\varphi^{-1}(p_{1},\ldots,p_{n})=(p_{1},\ldots,p_{n-1})^{\top}.$
+Let $\Delta$ be the standard simplex in $\mathbb{R}^{n}$, i.e.,
+$\Delta\coloneq\left\{(p_{i})\in\mathbb{R}^{n}\,\middle|\,\begin{gathered}p_{i}\geq 0\ (i=1,\ldots,n),\\ \textstyle\sum_{i=1}^{n}p_{i}=1\end{gathered}\right\}.$
 For $\lambda\in\mathbb{C}$, $\operatorname{Re}(\lambda)$, $\operatorname{Im}(\lambda)$, and $\overline{\lambda}$ denote the real part, the imaginary part, and the complex conjugate of $\lambda$, respectively.
@@
 For a real invertible matrix $X\in\mathbb{R}^{\ell\times\ell}$, we write $X^{-\top}\coloneq(X^{-1})^{\top}$, and for a complex invertible matrix $Y\in\mathbb{R}^{\ell\times\ell}$, we write $Y^{-*}\coloneq(Y^{-1})^{*}$.
-For a matrix $Z\in\mathbb{C}^{\ell\times\ell}$, $\mathrm{e}^{Z},\det Z$, and $\operatorname{tr}Z$ denote the exponential of $Z$, the determinant of $Z$, and the trace of $Z$.
-For $C=(c_{ij})\in\mathbb{R}^{\ell\times\ell}$ and $^{\prime}\in\mathbb{R}^{\ell^{\prime}\times\ell^{\prime}}, CC$$\otimes C^{\prime}\coloneq\begin{pmatrix}c_{11}C^{\prime}&\cdots&c_{1\ell}C^{\prime}\\ \vdots&\ddots&\vdots\\ c_{\ell 1}C^{\prime}&\cdots&c_{\ell\ell}C^{\prime}\end{pmatrix}\in\mathbb{R}^{\ell\ell^{\prime}\times\ell\ell^{\prime}}$ denotes their Kronecker product.
+For $Z\in\mathbb{C}^{\ell\times\ell}$, $\mathrm{e}^{Z},\det Z$, and $\operatorname{tr}Z$ denote the exponential, the determinant, and the trace of $Z$, respectively.
+For $C=(c_{ij})\in\mathbb{C}^{k\times\ell}$ and $^{\prime}\in\mathbb{C}^{k^{\prime}\times\ell^{\prime}}, CC$$\otimes C^{\prime}\coloneq\begin{pmatrix}c_{11}C^{\prime}&\cdots&c_{1\ell}C^{\prime}\\ \vdots&\ddots&\vdots\\ c_{k1}C^{\prime}&\cdots&c_{k\ell}C^{\prime}\end{pmatrix}\in\mathbb{C}^{kk^{\prime}\times\ell\ell^{\prime}}$ denotes their Kronecker product.
 For Hermitian matrices $X,Y\in\mathbb{C}^{\ell\times\ell}$, we write $X\succ Y$ if XYX-Y is positive definite.
 For $v_{1},\ldots,v_{\ell}\in\mathbb{C}$, let $\mathrm{diag}(v_{1},\ldots,v_{\ell})$ denote the diagonal matrix with the diagonal elements $v_{1},\ldots,v_{\ell}$.
-For a vector $v=(v_{1},\ldots,v_{\ell})^{\top}\in\mathbb{C}^{\ell}$, $\left\|{v}\right\|\coloneq\sqrt{\sum_{i=1}^{\ell}|v_{i}|^{2}}$ denotes the standard Euclidean norm.
-For a matrix $X\in\mathbb{C}^{k\times\ell}$, $\left\|{X}\right\|\coloneq\sqrt{\operatorname{tr}(X^{*}X)}$ denotes the Frobenius norm.
+For $v=(v_{1},\ldots,v_{\ell})^{\top}\in\mathbb{C}^{\ell}$, $\left\|{v}\right\|\coloneq\sqrt{\sum_{i=1}^{\ell}|v_{i}|^{2}}$ denotes the standard Euclidean norm, and for $X\in\mathbb{C}^{k\times\ell}$, $\left\|{X}\right\|\coloneq\sqrt{\operatorname{tr}(X^{*}X)}$ denotes the Frobenius norm.
 II Preliminaries
 II-A Controllability score
-In this subsection, we summarize the concept of the controllability scores [23].
-To define them, we introduce a virtual input allocation in which a fixed amount of control resources is distributed across network nodes.
-Let $p\in\Delta$ and consider the following system illustrated in Fig.
-1:
+We begin with an intuitive description of controllability scores [24] to motivate the formal definitions introduced later in this subsection.
+To this end, we consider the autonomous system (1) and introduce a virtual actuation mechanism (see Fig.
+1).
+Specifically, each state node $x_{i}$ is paired one-to-one with a virtual input $u_{i}$, and the input strength is determined by an allocation vector $p\in\Delta$:
 $\dot{x}(t)=Ax(t)+\operatorname{diag}(\sqrt{p_{1}},\ldots,\sqrt{p_{n}})u(t),$ (2)
-where $u(t)\coloneq(u_{1}(t),\ldots,u_{n}(t))^{\top}\in\mathbb{R}^{n}$ is a virtual control signal, and each node $x_{i}$ is actuated by $\sqrt{p_{i}}u_{i}(t)$.
+where $u(t)\coloneq(u_{1}(t),\ldots,u_{n}(t))^{\top}\in\mathbb{R}^{n}$.
 The value $p_{i}$ represents the amount of control resources allocated to node $x_{i}$; under a constraint on the input energy, a larger value of $p_{i}$ allows a stronger control effect on node $x_{i}$, indicating that node $x_{i}$ is a central node in the system.
-The controllability score associates each coefficient $p_{i}$ with the importance of node $x_{i}$.
+Given a controllability measure for system 2, we optimize it over $p$ and denote an optimal allocation by $p^{*}$.
+This optimal allocation $p^{*}$ maximizes controllability, and its entries $p_{i}^{*}$ quantify node importance.
+We therefore interpret $p^{*}$ as a dynamics-aware centrality measure, called a controllability score.
 $x_{1}$$x_{2}$$x_{3}$$x_{4}$$x_{5}$$u_{1}$$u_{2}$$u_{3}$$u_{4}$$u_{5}$$\sqrt{p_{1}}$$\sqrt{p_{2}}$$\sqrt{p_{3}}$$\sqrt{p_{4}}$$\sqrt{p_{5}}$
 Figure 1: Virtual system for defining controllability scores.
-The node $x_{i}$ denotes a state variable, and the node $u_{i}$ denotes a virtual control signal.
-The coefficients $p_{i}$ specify how a fixed amount of control resources is allocated to the nodes, and each input $u_{i}$ corresponds one-to-one with node $x_{i}$ through the weight $\sqrt{p_{i}}$.
-Let us maximize the controllability of system 2 with respect to $p$ for a certain measure, and let $p^{*}$ denote the optimal solution.
-The optimal solution $p^{*}$ represents the allocation of control authority that best enhances system controllability.
-The entries $p_{i}^{*}$ of this allocation, therefore, quantify how important each node $x_{i}$ is for controlling the system dynamics.
-Accordingly, $p^{*}$ can be regarded as a centrality measure, and we refer to it as a controllability score.
-The controllability of a system can be measured by the controllability Gramian. For $B\in\mathbb{C}^{n\times n}$, let $\widetilde{W}(B;T)\in\mathbb{C}^{n\times n}$ be defined as follows:
+Each node $x_{i}$ is paired one-to-one with a virtual input node $u_{i}$; the allocation $p$ distributes a fixed control budget, and $u_{i}$ acts on $x_{i}$ with gain $\sqrt{p_{i}}$.
+To turn this intuition into a formal definition, we introduce controllability-Gramian–based scalar measures of controllability as functions of $p$.
+We first define, for $B\in\mathbb{C}^{n\times n}$ and $T>0$,
 $\widetilde{W}(B;T)\coloneq\int_{0}^{T}\mathrm{e}^{At}B\mathrm{e}^{A^{\top}t}\mathrm{d}t.$ (3)
-The controllability Gramian for 2 with terminal time $T>0$ is given by $\widetilde{W}(\operatorname{diag}(p);T)$, which we abbreviate as $\widetilde{W}(p;T)$.
+The controllability Gramian for system 2 with terminal time $T>0$ is given by $\widetilde{W}(\operatorname{diag}(p);T)$, which we abbreviate as $\widetilde{W}(p;T)$.
 The quantities $\log\det\widetilde{W}(p;T)$ and $\operatorname{tr}\left(\widetilde{W}(p;T)^{-1}\right)$ can serve as scalar controllability measures.
@@
 Remark 1
-Note that centrality measures based on functions of the controllability Gramian are not necessarily well defined in the limit $T\to\infty$.
-For example, average controllability centrality (AC), VCE, and ACE proposed in [25] were originally defined for stable systems, but can be naively extended to unstable systems, in which case they are defined for node $i$ by
+Controllability-Gramian–based centralities need not have a meaningful limit as $T\to\infty$.
+For example, average controllability centrality (AC), VCE, and ACE proposed in [26] (originally for stable systems) can be formally extended to unstable systems by defining, for node $i$,
 $\mathrm{AC}_{i,T}\coloneq\operatorname{tr}\left(\widetilde{W}(e_{i}e_{i}^{\top};T)\right),$
@@
 where $k_{i}\coloneq\operatorname{rank}\left(\widetilde{W}(e_{i}e_{i}^{\top};T)\right)$ and $\lambda_{j}\left(\widetilde{W}(e_{i}e_{i}^{\top};T)\right)\ (j=1,\ldots,k_{i})$ denote the positive eigenvalues.
-Since eigenvalues of $\widetilde{W}(e_{i}e_{i}^{\top};T)$ that are close to zero reduce the values of $\mathrm{VCE}_{i,T}$ and $\mathrm{ACE}_{i,T}$, whereas zero eigenvalues do not, it has been pointed out in [23] that VCE and ACE do not necessarily provide appropriate centrality measures from the perspective of controllability.
-Furthermore, as discussed in [22], capturing the influence of network structure requires choosing the terminal time $T$ to be sufficiently large, which naturally leads one to consider the limit $T\to\infty$.
-Since $\left\|{\widetilde{W}(e_{i}e_{i}^{\top};T)}\right\|\to\infty$ as $T\to\infty$ for at least one $i$, the maximum eigenvalue of $\widetilde{W}(e_{i}e_{i}^{\top};T)$ also diverges, implying $\mathrm{AC}_{i,T}\to\infty$ and $\mathrm{VCE}_{i,T}\to\infty$.
+To capture the influence of network structure, one typically needs to take $T$ sufficiently large [23], which naturally motivates considering the limit $T\to\infty$.
+For unstable systems, however, one has $\left\|{\widetilde{W}(e_{i}e_{i}^{\top};T)}\right\|\to\infty$ as $T\to\infty$ for at least one $i$; hence the largest eigenvalue of $\widetilde{W}(e_{i}e_{i}^{\top};T)$ diverges, implying $\mathrm{AC}_{i,T}\to\infty$ and $\mathrm{VCE}_{i,T}\to\infty$.
+Even for finite $T$, VCE and ACE can be misleading as controllability-based centralities.
+In fact, as pointed out in [24], directions with very small (but positive) Gramian eigenvalues are heavily penalized, whereas directions with zero eigenvalues are omitted from the definitions and hence not penalized.
 In contrast, VCS and AECS are defined as optimal solutions to Problems 1 and 2, which is defined over a bounded feasible set.
@@
 III Scaling method of the controllability Gramian and reformulation of Problems 1 and 2
-In this section, we reformulate Problem 1 as
+To define the infinite-horizon controllability scores in Section IV, we first introduce a scaling of the controllability Gramian and derive numerically stable reformulations of the finite-horizon problems Problems 1 and 2.
+Specifically, in this section we reformulate Problem 1 as
 Problem 3
@@
 III-A Problem settings
-In this section, the matrix $A\in\mathbb{R}^{n\times n}$ is arbitrary, and we assume that it admits a Jordan decomposition
+Throughout this section, the matrix $A\in\mathbb{R}^{n\times n}$ is arbitrary, and we assume that it admits a Jordan decomposition
 $Q^{-1}AQ=J\coloneq\begin{pmatrix}J_{1}\\ &\ddots\\ &&J_{m}\end{pmatrix}$
@@
 Let $n_{-}\coloneq\sum_{i=1}^{m_{-}}n_{i},\ n_{0}\coloneq\sum_{i=m_{-}+1}^{m_{-}+m_{0}}n_{i}$, and $n_{+}\coloneq\sum_{i=m_{-}+m_{0}+1}^{m}n_{i}$ denote the numbers of eigenvalues with negative, zero, and positive real parts, respectively, counted with algebraic multiplicity.
-Let
-$J_{-}\coloneq\begin{pmatrix}J_{1}\\ &\ddots\\ &&J_{m_{-}}\end{pmatrix}\in\mathbb{C}^{n_{-}\times n_{-}}$
-be the Jordan matrix corresponding to the eigenvalues with negative real parts. Similarly, we define $J_{0}\in\mathbb{C}^{n_{0}\times n_{0}}$ and $J_{+}\in\mathbb{C}^{n_{+}\times n_{+}}$.
+Let $J_{-}\coloneq\begin{pmatrix}J_{1}\\ &\ddots\\ &&J_{m_{-}}\end{pmatrix}\in\mathbb{C}^{n_{-}\times n_{-}}$ be the Jordan matrix corresponding to the eigenvalues with negative real parts.
+Similarly, we define $J_{0}\in\mathbb{C}^{n_{0}\times n_{0}}$ and $J_{+}\in\mathbb{C}^{n_{+}\times n_{+}}$.
 Furthermore, let $m^{\prime}_{0}$ denote the number of distinct values among $\lambda_{m_{-}+1},\ldots,\lambda_{m_{-}+m_{0}}$, and let $\lambda^{\prime}_{1},\cdots,\lambda^{\prime}_{m^{\prime}_{0}}$ be these values, ordered so that $\operatorname{Im}(\lambda^{\prime}_{1})<\cdots<\operatorname{Im}(\lambda^{\prime}_{m^{\prime}_{0}})$.
@@
 Let $C_{1}\in\mathbb{C}^{\ell_{1}\times\ell_{1}},C_{2}\in\mathbb{C}^{\ell_{2}\times\ell_{2}}$, and $P\in\mathbb{C}^{\ell_{1}\times\ell_{2}}$.
-If $\mu_{1}+\mu_{2}\neq 0$ holds for any eigenvalues $\mu_{1}$ of $C_{1}$ and $\mu_{2}$ of $C_{2}$, then
-$C_{1}Y+YC_{2}+P=0$
-has a unique solution $Y\in\mathbb{C}^{\ell_{1}\times\ell_{2}}$.
+If $\mu_{1}+\mu_{2}\neq 0$ holds for any eigenvalues $\mu_{1}$ of $C_{1}$ and $\mu_{2}$ of $C_{2}$, then $C_{1}Y+YC_{2}+P=O$ has a unique solution $Y\in\mathbb{C}^{\ell_{1}\times\ell_{2}}$.
 Let us define $\widetilde{B}\coloneq Q^{-1}BQ^{-*}$ and partition $\widetilde{B}$ into blocks in the following two ways:
@@
 Then, let $W_{-}(B)$ and $W_{+}(B)$ be the unique solutions to
-$J_{-}W_{-}(B)+W_{-}(B)J_{-}^{*}+\widetilde{B}_{-,-}=0,$ (7)
-$(-J_{+})W_{+}(B)+W_{+}(B)(-J_{+}^{*})+\widetilde{B}_{+,+}=0.$ (8)
+$J_{-}W_{-}(B)+W_{-}(B)J_{-}^{*}+\widetilde{B}_{-,-}=O,$ (7)
+$(-J_{+})W_{+}(B)+W_{+}(B)(-J_{+}^{*})+\widetilde{B}_{+,+}=O.$ (8)
 For $i,j=m_{-}+1,\ldots,m_{-}+m_{0}$, we denote the $(n_{i},n_{j})$-entry of $\widetilde{B}_{i,j}$ by $\widetilde{b}_{i,j}$, which is the lower-right entry, and define
@@
 $W_{0}(B)\coloneq\begin{pmatrix}W_{0}^{(1)}(B)\\ &\ddots\\ &&W_{0}^{(m_{0}^{\prime})}(B)\end{pmatrix}$, (11)
-where we set $\alpha=m_{-}+m_{0}^{(1)}+\cdots+m_{0}^{(\ell-1)},\ \beta=m_{-}+m_{0}^{(1)}+\cdots+m_{0}^{(\ell)}$ to simplify the indexing.
+where we set $\alpha=m_{-}+m_{0}^{(1)}+\cdots+m_{0}^{(\ell-1)},\ \beta=\alpha+m_{0}^{(\ell)}$ to simplify the indexing.
 Then, the limit of $W(B;T)$ can be expressed in terms of $W_{-}(B),W_{0}(B)$, and $W_{+}(B)$.
@@
 Problem 4 is equivalent to Problem 2.
-Since the uniqueness of the optimal solutions is equivalent as well, we obtain the same result as in [23, Theorem 5] and [22, Theorem 1].
+Since the uniqueness of the optimal solutions is equivalent as well, we obtain the same result as in [24, Theorem 5] and [23, Theorem 1].
 In Appendix E, we summarize existing results on the uniqueness of the optimal solutions and present new results.
@@
 We then justify this formulation by proving that, under suitable assumptions, the optimal solutions are uniquely determined and that the finite-horizon controllability scores indeed converge to them.
-In what follows, we define
-$W_{\infty,i}\coloneq W_{\infty}(e_{i}e_{i}^{\top}),$
-for which
+In what follows, we define $W_{\infty,i}\coloneq W_{\infty}(e_{i}e_{i}^{\top})$, for which
 $W_{\infty}(p)=\sum_{i=1}^{n}p_{i}W_{\infty,i}$
@@
 $\mathrm{subject\ to}\quad$ $p\in\Delta,\ W_{\infty}(p)\succ O.$
+We refer to the optimal solution to Problem 5 as the VCS on an infinite time horizon, or simply the infinite-horizon VCS.
+Since $W(p;T)\to W_{\infty}(p)$ as $T\to\infty$, Problem 5 is obtained from Problem 3 by taking the limit $T\to\infty$, and its objective depends on the full matrix $W_{\infty}(p)$.
 Since the infinite-horizon version of Problem 4 is not well defined unless eigenvalues with negative real parts are present, we impose the following assumption.
@@
 $\mathrm{subject\ to}\quad$ $p\in\Delta,\ W_{-}(p)\succ O.$
-We refer to the optimal solution to Problem 5 as the VCS on an infinite time horizon, or simply the infinite-horizon VCS.
-Similarly, we refer to the optimal solution to Problem 6 as the AECS on an infinite time horizon, or simply the infinite-horizon AECS.
-In the case where $A$ is Hurwitz, Problems 5 and 6 coincide with the controllability scoring problems proposed in [23].
-Since $W(p;T)\to W_{\infty}(p)$, Problem 5 is formally obtained from Problem 3 by letting $T\to\infty$, and it depends on the whole matrix $W_{\infty}(p)$.
-On the other hand, since $D(T)^{-1}\to\begin{pmatrix}I_{n_{-}}\\ &O\end{pmatrix}$, formally letting $T\to\infty$ in Problem 4 yields the objective function of Problem 6, which depends only on $W_{-}(p)$; consequently, the constraint is posed only on $W_{-}(p)$.
+We refer to the optimal solution to Problem 6 as the AECS on an infinite time horizon, or simply the infinite-horizon AECS.
+As $T\to\infty$, we have $D(T)^{-1}\to\begin{pmatrix}I_{n_{-}}\\ &O\end{pmatrix}$, and Theorem 1 then shows that the objective in Problem 4 converges to that of Problem 6.
+In particular, the latter depends only on $W_{-}(p)$, and consequently the constraint is imposed only on $W_{-}(p)$.
+If $A$ is Hurwitz, then Problems 5 and 6 coincide with the controllability scoring problems proposed in [24].
 Remark 2
@@
 Consequently, while the infinite-horizon VCS enforces controllability of the full system, the infinite-horizon AECS may render the entire system uncontrollable.
+We illustrate this phenomenon with a numerical example in Section VI.
 IV-B Convergence of controllability scores
@@
 IV-C Uniqueness analysis
-We discuss the uniqueness of the optimal solutions to Problems 5 and 6. The uniqueness of the optimal solution is important for two reasons.
+We discuss uniqueness of the optimal solutions to Problems 5 and 6, which is important for two reasons.
 First, the convergence results for the finite-horizon controllability scores established in Section IV-B are derived under the assumption that the optimal solutions to Problems 5 and 6 are unique.
@@
 Indeed, by taking $Q=\dfrac{1}{\sqrt{2}}\begin{pmatrix}1&-1\\ 1&1\end{pmatrix}$, we obtain $Q^{-1}AQ=\operatorname{diag}(1,-1)$, whose eigenvalues are $1$ and 1-1.
-Therefore, by Theorem 7, Problems 1 and 2 each admit a unique optimal solution for all $T>0$.
+Therefore, $\Theta^{\prime}=\emptyset$ in Theorem 7 of Appendix E, and hence, by Theorem 7, Problems 1 and 2 each admit a unique optimal solution for all $T>0$.
 On the other hand, the infinite-horizon scaled controllability Gramians, are given by $W_{\infty,1}=W_{\infty,2}=\begin{pmatrix}\frac{1}{4}&0\\ 0&\frac{1}{4}\end{pmatrix}$ and $W_{-,1}=W_{-,2}=\frac{1}{4}$.
@@
 $\Box$
-We partition $Q^{-1}e_{i}$ into blocks in following two ways:
-$Q^{-1}e_{i}=\begin{pmatrix}q_{1,i}\\ \vdots\\ q_{m,i}\end{pmatrix}=\begin{pmatrix}q_{-,i}\\ q_{0,i}\\ q_{+,i}\end{pmatrix}$.
-Furthermore, let $q^{\prime}_{k,i}$ denote the $n_{k}$-th (i.e., the last) entry of $q_{k,i}$. For $\ell=1,\ldots,m_{0}^{\prime}$, we define
-$q_{0,i}^{(\ell)}\coloneq(q^{\prime}_{\alpha+1,i},\ldots,q^{\prime}_{\beta,i})^{\top}\in\mathbb{C}^{m_{0}^{(\ell)}},$
-where we set $\alpha=m_{-}+m_{0}^{(1)}+\cdots+m_{0}^{(\ell-1)},\ \beta=m_{-}+m_{0}^{(1)}+\cdots+m_{0}^{(\ell)}$ to simplify the indexing. Then, the following theorems hold.
+We partition $Q^{-1}e_{i}$ into blocks in following two ways: $Q^{-1}e_{i}=\begin{pmatrix}q_{1,i}\\ \vdots\\ q_{m,i}\end{pmatrix}=\begin{pmatrix}q_{-,i}\\ q_{0,i}\\ q_{+,i}\end{pmatrix}$.
+Furthermore, let $q^{\prime}_{k,i}$ denote the $n_{k}$-th (i.e., the last) entry of $q_{k,i}$.
+For $\ell=1,\ldots,m_{0}^{\prime}$, we define $q_{0,i}^{(\ell)}\coloneq(q^{\prime}_{\alpha+1,i},\ldots,q^{\prime}_{\beta,i})^{\top}\in\mathbb{C}^{m_{0}^{(\ell)}}$, where we set $\alpha=m_{-}+m_{0}^{(1)}+\cdots+m_{0}^{(\ell-1)},\ \beta=\alpha+m_{0}^{(\ell)}$ to simplify the indexing.
+Then, the following theorems hold.
 Theorem 4
@@
 V Algorithms
-In this section, we describe an algorithm for solving Problems 5 and 6.
+As discussed in Section I, the observations in [23] suggest choosing $T$ sufficiently large in Problems 1 and 2, in which case the resulting problems can be solved using existing algorithms [24].
+On the other hand, Theorems 2 and 3 show that Problems 5 and 6 yield the limiting values of these finite-horizon controllability scores.
+Moreover, beyond yielding the limit, Problems 5 and 6 admit a numerically stable computation—even for unstable systems—via the scaled-Gramian reformulation, whereas the direct finite-horizon computation can suffer from numerical instability when $T$ is large.
+Finally, solving Problems 5 and 6 is computationally more efficient than solving Problems 1 and 2 for large values of $T$, as described below.
 In formulating Problems 5 and 6, we employed the Jordan canonical form; however, a direct use of the Jordan canonical form may lead to numerical instability [8].
-As shown below, under the following assumption, the optimal solutions to Problems 5 and 6 can be computed in a numerically stable manner via the real Schur decomposition and the solutions of Sylvester equations.
+As shown below, under the following assumption, the optimal solutions to Problems 5 and 6 can be computed in a numerically stable manner via the real Schur decomposition and solutions of Sylvester equations.
 Assumption 2
@@
 […]
-and the diagonal entries satisfy
-$\ell_{ii}=-\sum_{j\neq i}\ell_{ij}.$
+and the diagonal entries satisfy $\ell_{ii}=-\sum_{j\neq i}\ell_{ij}$.
 When considering graph Laplacians, we restrict attention to positive edge weights whenever an edge exists.
 For undirected graphs, the Laplacian matrix $L$ is symmetric, whereas for directed graphs, $L$ is generally not symmetric.
-Let $L$ be the Laplacian matrix of a directed graph. Then, A=LA=-L satisfies Assumption 2. In fact, from the definition, the matrix $A=(a_{ij})$ satisfies
-$a_{ij}\geq 0\ (i\neq j),\quad a_{ii}=-\sum_{j\neq i}a_{ij}\leq 0.$
-It follows from Geršhgorin’s theorem [27, Theorem 1.1] that all eigenvalues of $A$ have nonpositive real parts, and that zero is the only eigenvalue that may lie on the imaginary axis.
+Let $L$ be the Laplacian matrix of a directed graph.
+Then, A=LA=-L satisfies Assumption 2.
+In fact, from the definition, the matrix $A=(a_{ij})$ satisfies $a_{ij}\geq 0\ (i\neq j)$ and $a_{ii}=-\sum_{j\neq i}a_{ij}\leq 0$.
+It follows from Geršhgorin’s theorem [28, Theorem 1.1] that all eigenvalues of $A$ have nonpositive real parts, and that zero is the only eigenvalue that may lie on the imaginary axis.
 Furthermore, $\operatorname{rank}(A)=\operatorname{rank}(A^{2})$ holds [6, Proposition 12].
@@
 $\Box$
-First, we perform a block diagonalization using the real Schur decomposition and the solution of a Sylvester equation.
+First, we perform a block diagonalization using the real Schur decomposition and the solution of Sylvester equations.
 Since this approach does not require the Jordan canonical form, it is numerically stable.
@@
 $A_{+}R^{(2)}_{+,-}-R^{(2)}_{+,-}A_{-}=-(A_{+,-}+A_{+,0}R^{(2)}_{0,-})$. (16)
-Since no two of the matrices $A_{-},A_{0}$, and $A_{+}$ share a common eigenvalue, 14, 15, and 16 can be solved successively and admit a unique solution from Lemma 1. Define
-$R^{(2)}\coloneq\begin{pmatrix}I_{n_{-}}\\ R^{(2)}_{0,-}&I_{n_{0}}\\ R^{(2)}_{+,-}&R^{(2)}_{+,0}&I_{n_{+}}\end{pmatrix}$
-and set $R=R^{(1)}R^{(2)}$. Then, by using
+Since no two of the matrices $A_{-},A_{0}$, and $A_{+}$ share a common eigenvalue, 14, 15, and 16 can be solved successively and admit a unique solution from Lemma 1.
+Define $R^{(2)}\coloneq\begin{pmatrix}I_{n_{-}}\\ R^{(2)}_{0,-}&I_{n_{0}}\\ R^{(2)}_{+,-}&R^{(2)}_{+,0}&I_{n_{+}}\end{pmatrix}$ and set $R=R^{(1)}R^{(2)}$.
+Then, by using
 $\left(R^{(2)}\right)^{-1}=\begin{pmatrix}I_{n_{-}}\\ -R^{(2)}_{0,-}&I_{n_{0}}\\ -R^{(2)}_{+,-}+R^{(2)}_{+,0}R^{(2)}_{0,-}&-R^{(2)}_{+,0}&I_{n_{+}}\end{pmatrix}$, (17)
@@
 where $r_{-,i}\in\mathbb{R}^{n_{-}},\ r_{0,i}\in\mathbb{R}^{n_{0}},\ r_{+,i}\in\mathbb{R}^{n_{+}}$, and let $V_{-,i}\in\mathbb{R}^{n_{-}\times n_{-}}$ and $V_{+,i}\in\mathbb{R}^{n_{+}\times n_{+}}$ be the solutions to
-$A_{-}V_{-,i}+V_{-,i}A_{-}^{\top}+q_{-,i}r_{-,i}^{\top}=0,$ (20)
-$A_{+}V_{+,i}+V_{+,i}A_{+}^{\top}-q_{+,i}r_{+,i}^{\top}=0.$ (21)
-Moreover, let $V_{0,i}\coloneq r_{0,i}r_{0,i}^{\top}$. Finally, for $p=(p_{1},\ldots,p_{n})^{\top}\in\mathbb{R}^{n}$, we define
+$A_{-}V_{-,i}+V_{-,i}A_{-}^{\top}+q_{-,i}r_{-,i}^{\top}=O,$ (20)
+$(-A_{+})V_{+,i}+V_{+,i}(-A_{+}^{\top})+q_{+,i}r_{+,i}^{\top}=O.$ (21)
+Moreover, let
+$V_{0,i}\coloneq r_{0,i}r_{0,i}^{\top}.$ (22)
+Finally, for $p=(p_{1},\ldots,p_{n})^{\top}\in\mathbb{R}^{n}$, we define
 $V_{\alpha}(p)\coloneq\sum_{i=1}^{n}p_{i}V_{\alpha,i}\quad(\alpha\in\{-,0,+\}),$
@@
 Thus, there exist invertible matrices $S_{-}\in\mathbb{C}^{n_{-}\times n_{-}},\ S_{0}\in\mathbb{C}^{n_{0}\times n_{0}}$, and $S_{+}\in\mathbb{C}^{n_{+}\times n_{+}}$ such that $S_{-}^{-1}A_{-}S_{-}=J_{-},\ S_{0}^{-1}A_{0}S_{0}=J_{0}$, and $S_{+}^{-1}A_{+}S_{+}=J_{+}$.
-Moreover, letting
-$S\coloneq\begin{pmatrix}S_{-}\\ &S_{0}\\ &&S_{+}\end{pmatrix}$,
-we obtain $S^{-1}(R^{-1}AR)S=J$; hence $S_{-},S_{0}$ and $S_{+}$ can be chosen so that $RS=Q$.
-As defined in Section IV-C,
-$\begin{pmatrix}q_{-,i}\\ q_{0,i}\\ q_{+,i}\end{pmatrix}=Q^{-1}e_{i}=S^{-1}\begin{pmatrix}r_{-,i}\\ r_{0,i}\\ r_{+,i}\end{pmatrix}$.
+Moreover, letting $S\coloneq\begin{pmatrix}S_{-}\\ &S_{0}\\ &&S_{+}\end{pmatrix}$, we obtain $S^{-1}(R^{-1}AR)S=J$; hence $S_{-},S_{0}$ and $S_{+}$ can be chosen so that $RS=Q$.
+As defined in Section IV-C, $\begin{pmatrix}q_{-,i}\\ q_{0,i}\\ q_{+,i}\end{pmatrix}=Q^{-1}e_{i}=S^{-1}\begin{pmatrix}r_{-,i}\\ r_{0,i}\\ r_{+,i}\end{pmatrix}$.
 Then, $r_{-,i}=S_{-}q_{-,i},\ r_{0,i}=S_{0}q_{0,i}$, and $r_{+,i}=S_{+}q_{+,i}$ hold.
 From 7 and 8, $W_{-,i}$ and $W_{+,i}$ are the unique solutions to
-$J_{-}W_{-,i}+W_{-,i}J_{-}^{*}+r_{-,i}r_{-,i}^{*}=0,$ (22)
-$(-J_{+})W_{+,i}+W_{+,i}(-J_{+}^{*})+r_{+,i}r_{+,i}^{*}=0,$ (23)
+$J_{-}W_{-,i}+W_{-,i}J_{-}^{*}+r_{-,i}r_{-,i}^{*}=O,$ (23)
+$(-J_{+})W_{+,i}+W_{+,i}(-J_{+}^{*})+r_{+,i}r_{+,i}^{*}=O,$ (24)
 respectively.
-Multiplying 22 from the left by $S_{-}$ and from the right by $S_{-}^{*}$, and using the uniqueness of solutions to 20, we obtain $V_{-,i}=S_{-}W_{-,i}S_{-}^{*}$.
-Similary, it follows from 21 and 23 that $V_{+,i}=S_{+}W_{+,i}S_{+}^{*}$.
+Multiplying 23 from the left by $S_{-}$ and from the right by $S_{-}^{*}$, and using the uniqueness of solutions to 20, we obtain $V_{-,i}=S_{-}W_{-,i}S_{-}^{*}$.
+Similary, it follows from 21 and 24 that $V_{+,i}=S_{+}W_{+,i}S_{+}^{*}$.
 Since $W_{0,i}$ reduces to $q_{0,i}q_{0,i}^{*}$ under Assumption 2, we obtain $V_{0,i}=S_{0}W_{0,i}S_{0}^{*}$.
@@
 $=\operatorname{tr}\left(R^{(1)}(R^{(2)})^{-\top}\begin{pmatrix}V_{-}(p)^{-1}\\ &O\end{pmatrix}(R^{(2)})^{-1}(R^{(1)})^{\top}\right)$
-$=\operatorname{tr}\left(R^{(1)}\begin{pmatrix}V_{-}(p)^{-1}\\ &O\end{pmatrix}(R^{(1)})^{\top}\right)$
-$=\operatorname{tr}\left(V_{-}(p)^{-1}\right),$
+$=\operatorname{tr}\left(R^{(1)}\begin{pmatrix}V_{-}(p)^{-1}\\ &O\end{pmatrix}(R^{(1)})^{\top}\right)=\operatorname{tr}\left(V_{-}(p)^{-1}\right),$
 where we used the fact that $R^{(1)}$ is orthogonal and that $R^{(2)}$ is lower triangular.
@@
 Algorithm 1 The computation of $V_{-,i},V_{0,i}$, and $V_{+,i}$
-Once $V_{\mathrm{l},i},V_{\mathrm{i},i}$, and $V_{\mathrm{r},i}$ have been computed, Problems 7 and 8 take the same form as the Problems 1 and 2.
+Once $V_{-,i},V_{0,i}$, and $V_{+,i}$ have been computed, Problems 7 and 8 take the same form as the Problems 1 and 2.
 Hence, we can employ Algorithms 2 and 3, which are the same as those in [24].
@@
 Let the objective functions of Problems 7 and 8 be
-$g(p)\coloneq-\log\det V(p),$
-$h(p)\coloneq\operatorname{tr}\left(V_{-}(p)^{-1}\right).$
+$g(p)\coloneq-\log\det V(p),\quad h(p)\coloneq\operatorname{tr}\left(V_{-}(p)^{-1}\right).$
 Then, their gradients can be computed by
@@
 The same argument applies to the AECS case.
+When $T$ is large but finite, computing $\widetilde{W}(e_{i}e_{i}^{\top};T)$ or $W(e_{i}e_{i}^{\top};T)$ via numerical integration leads to a high computational cost that increases with $T$.
+In contrast, matrices $V_{i}$ that play the same role as the infinite-horizon scaled controllability Gramians $W_{\infty,i}$ can be computed by solving the Lyapunov equations 20 and 21 together with computing 22, whose computational costs are independent of $T$.
+Moreover, Lyapunov equations can be solved efficiently in an approximate manner using algorithms such as CF-ADI [15], making the infinite-horizon formulation advantageous from a computational perspective.
 VI Numerical experiments
@@
 Under suitable assumptions, we proved that the finite-horizon controllability scores converge to the infinite-horizon ones and that the infinite-horizon controllability scores are unique.
-This proves a principled choice of the terminal time, namely $T=\infty$.
+This provides a principled choice of the terminal time, namely $T=\infty$.
 Although the formulations of infinite-horizon controllability scoring problems rely on the Jordan canonical form of $A$, we also showed that, under appropriate assumptions, their optimal solutions can be computed using the Schur decomposition and the solution of Sylvester equations, and we presented a corresponding numerical algorithm.
+Moreover, the infinite-horizon formulation is more advantageous than the finite-horizon one from a computational perspective.
 In the numerical experiments, we computed infinite-horizon controllability scores for the same Laplacian dynamics as in [23] and observed that finite-horizon controllability scores indeed converge to infinite-horizon ones as $T\to\infty$.
@@
 Lemma 4 ([14, Theorem 5])
-Let $C_{1}\in\mathbb{C}^{\ell_{1}\times\ell_{1}},C_{2}\in\mathbb{C}^{\ell_{2}\times\ell_{2}}$, and $R\in\mathbb{C}^{\ell_{1}\times\ell_{2}}$.
-If $\operatorname{Re}(\mu_{1}+\mu_{2})<0$ holds for any eigenvalues $\mu_{1}$ of $C_{1}$ and $\mu_{2}$ of $C_{2}$, then
-$\int_{0}^{\infty}\mathrm{e}^{C_{1}t}P\mathrm{e}^{C_{2}t}\mathrm{d}t$
-converges, and it is the unique solution $Y\in\mathbb{C}^{\ell_{1}\times\ell_{2}}$ satisfying
-$C_{1}Y+YC_{2}+P=0.$
+Let $C_{1}\in\mathbb{C}^{\ell_{1}\times\ell_{1}},C_{2}\in\mathbb{C}^{\ell_{2}\times\ell_{2}}$, and $P\in\mathbb{C}^{\ell_{1}\times\ell_{2}}$.
+If $\operatorname{Re}(\mu_{1}+\mu_{2})<0$ holds for any eigenvalues $\mu_{1}$ of $C_{1}$ and $\mu_{2}$ of $C_{2}$, then $\int_{0}^{\infty}\mathrm{e}^{C_{1}t}P\mathrm{e}^{C_{2}t}\mathrm{d}t$ converges, and it is the unique solution $Y\in\mathbb{C}^{\ell_{1}\times\ell_{2}}$ satisfying $C_{1}Y+YC_{2}+P=0$.
 Proof (Theorem 1)
@@
 $\mathrm{e}^{Jt}\widetilde{B}\mathrm{e}^{J^{*}t}=\begin{pmatrix}\mathrm{e}^{J_{1}t}\widetilde{B}_{1,1}\mathrm{e}^{J_{1}^{*}t}&\cdots&\mathrm{e}^{J_{1}t}\widetilde{B}_{1,m}\mathrm{e}^{J_{m}^{*}t}\\ \vdots&\ddots&\vdots\\ \mathrm{e}^{J_{m}t}\widetilde{B}_{m,1}\mathrm{e}^{J_{1}^{*}t}&\cdots&\mathrm{e}^{J_{m}t}\widetilde{B}_{m,m}\mathrm{e}^{J_{m}^{*}t}\end{pmatrix}$ (28)
-From 4, 5, 26, and 27, the $(i,j)$-block of $W(B;T)$ can be written as
+From 4, 5, 27, and 28, the $(i,j)$-block of $W(B;T)$ can be written as
 $D_{i}(T)^{-1}\left\{\int_{0}^{T}\mathrm{e}^{J_{i}t}\widetilde{B}_{i,j}\mathrm{e}^{J_{j}^{*}t}\mathrm{d}t\right\}D_{j}(T)^{-*}.$ (29)
-In the following, we analyze the block representation 28 by distinguishing the following nine cases:
+In the following, we analyze the block representation 29 by distinguishing the following nine cases:
   1. 1)
@@
 Case 1)
-Noting that $D_{i}(T)=I_{n_{i}}$ and $D_{j}(T)=I_{n_{j}}$, 28 reduces to
-$\int_{0}^{T}\mathrm{e}^{J_{i}t}\widetilde{B}_{i,j}\mathrm{e}^{J_{j}^{*}t}\mathrm{d}t.$
+Noting that $D_{i}(T)=I_{n_{i}}$ and $D_{j}(T)=I_{n_{j}}$, 29 reduces to $\int_{0}^{T}\mathrm{e}^{J_{i}t}\widetilde{B}_{i,j}\mathrm{e}^{J_{j}^{*}t}\mathrm{d}t$.
 Since $\operatorname{Re}(\lambda_{i}),\operatorname{Re}(\lambda_{j})<0$, it follows from Lemma 4 that the $(i,j)$-block converges, as $T\to\infty$, to the unique solution $W_{-}^{(i,j)}(B)\in\mathbb{C}^{n_{i}\times n_{j}}$ satisfying
-$J_{i}W_{-}^{(i,j)}(B)+W_{-}^{(i,j)}(B)J_{j}^{*}+\widetilde{B}_{i,j}=0.$
+$J_{i}W_{-}^{(i,j)}(B)+W_{-}^{(i,j)}(B)J_{j}^{*}+\widetilde{B}_{i,j}=O.$
 Then, the upper-left block of $W(B;T)$ of size $n_{-}\times n_{-}$ converges to
@@
 Case 2)
-Noting that $D_{i}(T)=I_{n_{i}}$, 28 reduces to
+Noting that $D_{i}(T)=I_{n_{i}}$, 29 reduces to
 $\left\{\int_{0}^{T}\mathrm{e}^{J_{i}t}\widetilde{B}_{i,j}\mathrm{e}^{J_{j}^{*}t}\mathrm{d}t\right\}D_{j}(T)^{-*}.$
@@
 Case 3)
-Noting that $D_{i}(T)=I_{n_{i}}$ and $D_{j}(T)=\mathrm{e}^{J_{j}T}$, 28 reduces to
+Noting that $D_{i}(T)=I_{n_{i}}$ and $D_{j}(T)=\mathrm{e}^{J_{j}T}$, 29 reduces to
 $\left\{\int_{0}^{T}\mathrm{e}^{J_{i}t}\widetilde{B}_{i,j}\mathrm{e}^{J_{j}^{*}t}\mathrm{d}t\right\}\mathrm{e}^{-J_{j}^{*}T}.$
-There exists a constant $c\geq 0$ such that
-$\left\|{\mathrm{e}^{J_{i}t}}\right\|\leq ct^{n_{i}-1}\mathrm{e}^{\operatorname{Re}(\lambda_{i})t},$
-$\left\|{\mathrm{e}^{J_{j}^{*}t}}\right\|\leq ct^{n_{j}-1}\mathrm{e}^{\operatorname{Re}(\lambda_{j})t},$
-$\left\|{\widetilde{B}_{i,j}}\right\|\leq c$
-for all $t\geq 0$. Consequently, we obtain
+There exists a constant $c\geq 0$ such that $\left\|{\mathrm{e}^{J_{i}t}}\right\|\leq ct^{n_{i}-1}\mathrm{e}^{\operatorname{Re}(\lambda_{i})t}$, $\left\|{\mathrm{e}^{J_{j}^{*}t}}\right\|\leq ct^{n_{j}-1}\mathrm{e}^{\operatorname{Re}(\lambda_{j})t}$, and $\left\|{\widetilde{B}_{i,j}}\right\|\leq c$ for all $t\geq 0$.
+Consequently, we obtain
 $\left\|{\int_{0}^{T}\mathrm{e}^{J_{i}t}\widetilde{B}_{i,j}\mathrm{e}^{J_{j}^{*}t}\mathrm{d}t}\right\|$
@@
 Case 5)
-Since
-$\mathrm{e}^{J_{i}t}=\sum_{k=0}^{n_{i}-1}\mathrm{e}^{\lambda_{i}t}\dfrac{t^{k}}{k!}N_{n_{i}}^{k},$
-we obtain
+Since $\mathrm{e}^{J_{i}t}=\sum_{k=0}^{n_{i}-1}\mathrm{e}^{\lambda_{i}t}\dfrac{t^{k}}{k!}N_{n_{i}}^{k}$, we obtain
 $\mathrm{e}^{J_{i}t}\widetilde{B}_{i,j}\mathrm{e}^{J_{j}^{*}t}=\sum_{k=0}^{n_{i}-1}\sum_{\ell=0}^{n_{j}-1}\mathrm{e}^{(\lambda_{i}+\overline{\lambda_{j}})t}\dfrac{t^{k+\ell}}{k!\ell!}N_{n_{i}}^{k}\widetilde{B}_{i,j}(N_{n_{j}}^{*})^{\ell}.$
-Consequently, 28 can be written as
+Consequently, 29 can be written as
 $\sum_{k=0}^{n_{i}-1}\sum_{\ell=0}^{n_{j}-1}\left\{\int_{0}^{T}\mathrm{e}^{(\lambda_{i}+\overline{\lambda_{j}})t}\dfrac{t^{k+\ell}}{k!\ell!}\mathrm{d}t\right\}\times\\ D_{i}(T)^{-1}N_{n_{i}}^{k}\widetilde{B}_{i,j}(N_{n_{j}}^{*})^{\ell}D_{j}(T)^{-*}.$ (30)
@@
 $\left\{\begin{aligned} \left|\int_{0}^{T}\mathrm{e}^{(\lambda_{i}+\overline{\lambda_{j}})t}\dfrac{t^{k+\ell}}{k!\ell!}\mathrm{d}t\right|&\leq cT^{k+\ell}&\text{if $\lambda_{i}\neq\lambda_{j}$},\\ \int_{0}^{T}\mathrm{e}^{(\lambda_{i}+\overline{\lambda_{j}})t}\dfrac{t^{k+\ell}}{k!\ell!}\mathrm{d}t&=\dfrac{T^{k+\ell+1}}{k!\ell!(k+\ell+1)}&\text{if $\lambda_{i}=\lambda_{j}$}\end{aligned}\right$.
-for some constant $c\geq 0$. Consequently, the summand of 29 converges, as $T\to\infty$, to
+for some constant $c\geq 0$. Consequently, the summand of 30 converges, as $T\to\infty$, to
 $\left\{\begin{aligned} &0&\text{if $\lambda_{i}\neq\lambda_{j}$},\\ &\dfrac{1}{k!\ell!(k+\ell+1)}E_{i,k}N_{n_{i}}^{k}\widetilde{B}_{i,j}(N_{n_{j}}^{*})^{\ell}E_{j,\ell}^{*}&\text{if $\lambda_{i}=\lambda_{j}$}.\end{aligned}\right$.
@@
 Case 6)
-Noting that $D_{j}(T)=\mathrm{e}^{J_{j}T}$, we can rewrite 28 as
+Noting that $D_{j}(T)=\mathrm{e}^{J_{j}T}$, we can rewrite 29 as
 $D_{i}(T)^{-1}\mathrm{e}^{J_{i}T}\left\{\int_{0}^{T}\mathrm{e}^{J_{i}(t-T)}\widetilde{B}_{i,j}\mathrm{e}^{J_{j}^{*}(t-T)}\mathrm{d}t\right\}$
@@
 Case 9)
-Noting that $D_{i}(T)=\mathrm{e}^{J_{i}T}$ and $D_{j}(T)=\mathrm{e}^{J_{j}T}$, 28 reduces to
-$\int_{0}^{T}\mathrm{e}^{J_{i}(t-T)}\widetilde{B}_{i,j}\mathrm{e}^{J_{j}^{*}(t-T)}\mathrm{d}t$
-$=\int_{0}^{T}\mathrm{e}^{-J_{i}t}\widetilde{B}_{i,j}\mathrm{e}^{-J_{j}^{*}t}\mathrm{d}t.$
+Noting that $D_{i}(T)=\mathrm{e}^{J_{i}T}$ and $D_{j}(T)=\mathrm{e}^{J_{j}T}$, 29 reduces to $\int_{0}^{T}\mathrm{e}^{J_{i}(t-T)}\widetilde{B}_{i,j}\mathrm{e}^{J_{j}^{*}(t-T)}\mathrm{d}t=\int_{0}^{T}\mathrm{e}^{-J_{i}t}\widetilde{B}_{i,j}\mathrm{e}^{-J_{j}^{*}t}\mathrm{d}t$.
 Since $\operatorname{Re}(\lambda_{i}),\operatorname{Re}(\lambda_{j})>0$, it follows from Lemma 4 that the integral of the right-hand side converges, as $T\to\infty$, to the unique solution $W_{+}^{(i,j)}(B)\in\mathbb{C}^{n_{i}\times n_{j}}$ satisfying
@@
 Therefore, since $\operatorname{diag}(p)$ is positive definite,
-$y^{\top}\widetilde{W}(p;T)y$ $=\int_{0}^{T}y^{\top}\mathrm{e}^{At}\operatorname{diag}(p)\mathrm{e}^{A^{\top}t}y\mathrm{d}t$
-$>0.$
+$y^{\top}\widetilde{W}(p;T)y=\int_{0}^{T}y^{\top}\mathrm{e}^{At}\operatorname{diag}(p)\mathrm{e}^{A^{\top}t}y\mathrm{d}t>0.$
 This proves that $\widetilde{W}(p;T)$ is positive definite. $\Box$
@@
 Lemma 5 ([14, Theorem 12])
-Let $C\in\mathbb{C}^{\ell\times\ell}$ be Hurwitz and $P\in\mathbb{C}^{\ell\times\ell}$ be Hermitian positive definite. Then, the Lyapunov equation
-$CY+YC^{*}+P=0$
-admits a unique solution $Y\in\mathbb{C}^{\ell\times\ell}$, which is Hermitian positive definite.
+Let $C\in\mathbb{C}^{\ell\times\ell}$ be Hurwitz and $P\in\mathbb{C}^{\ell\times\ell}$ be Hermitian positive definite.
+Then, the Lyapunov equation $CY+YC^{*}+P=O$ admits a unique solution $Y\in\mathbb{C}^{\ell\times\ell}$, which is Hermitian positive definite.
 Proof (Lemma 3)
@@
 From 7, the matrix $W_{-}(p)$ is the unique solution to
-$J_{-}W_{-}(p)+W_{-}(p)J_{-}^{*}+\widetilde{B}_{-,-}=0,$
+$J_{-}W_{-}(p)+W_{-}(p)J_{-}^{*}+\widetilde{B}_{-,-}=O,$
 and it follows from Lemma 5 that $W_{-}(p)$ is positive definite. Similarly, we can prove that $W_{+}(p)$ is positive definite.
@@
 Since $W_{0}(p)$ is a block-diagonal matrix from 11, it suffices to show that $W_{0}^{(\ell)}(p)$ is positive definite for $\ell=1,\ldots,m_{0}^{\prime}$.
-To simplify the indexing, let $\alpha=m_{-}+m_{0}^{(1)}+\cdots+m_{0}^{(\ell-1)},\ \beta=m_{-}+m_{0}^{(1)}+\cdots+m_{0}^{(\ell)}$.
+To simplify the indexing, let $\alpha=m_{-}+m_{0}^{(1)}+\cdots+m_{0}^{(\ell-1)},\ \beta=\alpha+m_{0}^{(\ell)}$.
 Let $e^{(i)}\coloneq(0,\ldots,0,1)^{\top}\in\mathbb{R}^{n_{i}\times n_{i}}$.
@@
 which is positive definite since it is a principal submatrix of $\widetilde{B}_{0,0}$,
-$E^{(\ell)}$ $\coloneq\begin{pmatrix}e^{(\alpha+1)}&0&\cdots&0&0\\ 0&e^{(\alpha+2)}&\cdots&0&0\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&\cdots&e^{(\beta-1)}&0\\ 0&0&\cdots&0&e^{(\beta)}\end{pmatrix}$
-$\in\mathbb{R}^{n_{0}^{(\ell)}\times m_{0}^{(\ell)}},$
+$E^{(\ell)}\coloneq\begin{pmatrix}e^{(\alpha+1)}&\cdots&0\\ \vdots&\ddots&\vdots\\ 0&\cdots&e^{(\beta)}\end{pmatrix}\in\mathbb{R}^{n_{0}^{(\ell)}\times m_{0}^{(\ell)}},$
 and
@@
 Appendix C Proof of Theorems 2 and 3
+Before proving Theorems 2 and 3, we collect several definitions and preliminary facts that will be used throughout the proofs.
+Let $\widetilde{\Delta}$ be the simplex in $\mathbb{R}^{n-1}$, i.e.,
+$\widetilde{\Delta}\coloneq\left\{(\widetilde{p}_{i})\in\mathbb{R}^{n-1}\,\middle|\,\begin{gathered}\widetilde{p}_{i}\geq 0\ (i=1,\ldots,n-1),\\ \textstyle\sum_{i=1}^{n-1}\widetilde{p}_{i}\leq 1\end{gathered}\right\}.$
+The mapping $\varphi\colon\widetilde{\Delta}\to\Delta$ defined by
+$\varphi(\widetilde{p}_{1},\ldots,\widetilde{p}_{n-1})=(\widetilde{p}_{1},\ldots,\widetilde{p}_{n-1},1-\sum_{i=1}^{n-1}\widetilde{p}_{i})^{\top}$
+is a homeomorphism; its inverse is given by
+$\varphi^{-1}(p_{1},\ldots,p_{n})=(p_{1},\ldots,p_{n-1})^{\top}.$
+For a set $\Omega\subset\mathbb{R}^{n-1}$, $\operatorname*{int}\Omega$ denotes its interior.
 […]
 Let $\{f^{(k)}\}_{k\in\mathbb{N}}$ be a sequence of functions $f^{(k)}\colon\mathbb{R}^{n-1}\to\mathbb{R}\cup\{\infty\}$.
-We say that $\{f^{(k)}\}_{k\in\mathbb{N}}$ is equi level-bounded if there exists a bounded set $B\subset\mathbb{R}^{n-1}$ such that $\mathrm{lev}_{\leq\alpha}f^{(k)}\subset B$ for all $k$.
-We say that $\{f^{(k)}\}_{k\in\mathbb{N}}$ epi-converges to $f$ if, for every $p^{\prime}\in\mathbb{R}^{n-1}$,
+We say that $\{f^{(k)}\}_{k\in\mathbb{N}}$ is equi level-bounded if for every $\alpha$, there exists a bounded set $B\subset\mathbb{R}^{n-1}$ such that $\mathrm{lev}_{\leq\alpha}f^{(k)}\subset B$ for all $k$.
+We say that $\{f^{(k)}\}_{k\in\mathbb{N}}$ epi-converges to $f$ if, for every $\widetilde{p}\in\mathbb{R}^{n-1}$,
 […]
 We employ the following three propositions.
-Proposition 1 ([21, Theorem 1.9])
+Proposition 1 ([22, Theorem 1.9])
 Assume that $f\colon\mathbb{R}^{n-1}\to\mathbb{R}\cup\{\infty\}$ is proper, lower semicontinuous, and level-bounded.
 Then, $\operatorname*{arg\,min}f$ is nonempty and compact, and the minimum value of $f$ is finite.
-Proposition 2 ([21, Theorem 7.17])
+Proposition 2 ([22, Theorem 7.17])
 […]
-Proposition 3 ([21, Theorem 7.33])
+Proposition 3 ([22, Theorem 7.33])
 Assume that $\{f^{(k)}\}_{k\in\mathbb{N}}$ is an equi level-bounded sequence of functions $f^{(k)}\colon\mathbb{R}^{n-1}\to\mathbb{R}\cup\{\infty\}$, that $\{f^{(k)}\}_{k\in\mathbb{N}}$ epi-converges to $f$, and that $f^{(k)}$ and $f$ are lower semicontinuous and proper.
-Then, the sequence $\{\widetilde{p}^{(k)}\}_{k\in\mathbb{N}}$ such that $p^{(k)}\in\operatorname*{arg\,min}f^{(k)}$ is bounded, and all its cluster points lie in $\operatorname*{arg\,min}f$.
+Then, the sequence $\{\widetilde{p}^{(k)}\}_{k\in\mathbb{N}}$ such that $\widetilde{p}^{(k)}\in\operatorname*{arg\,min}f^{(k)}$ is bounded, and all its cluster points lie in $\operatorname*{arg\,min}f$.
 In particular, when $\operatorname*{arg\,min}f=\{\widetilde{p}^{*}\}$, the sequence $\{\widetilde{p}^{(k)}\}_{k\in\mathbb{N}}$ converges to $\widetilde{p}^{*}$.
@@
 Lemma 6
-Let $C_{1},C_{2}\in\mathbb{C}^{\ell\times\ell}$ be Hermitian positive definite matrices, and $\mu>0$ is the minimum eigenvalue of $C_{2}$. Then,
-$\operatorname{tr}(C_{1}C_{2})\geq\mu\operatorname{tr}(C_{1})$
-holds.
+Let $C_{1},C_{2}\in\mathbb{C}^{\ell\times\ell}$ be Hermitian positive definite matrices, and $\mu>0$ is the minimum eigenvalue of $C_{2}$.
+Then, $\operatorname{tr}(C_{1}C_{2})\geq\mu\operatorname{tr}(C_{1})$ holds.
 Proof
-Since $C_{2}\succeq\mu I_{\ell}$, we obtain $C_{1}^{1/2}C_{2}C_{1}^{1/2}\succeq\mu C_{1}$. Therefore,
-$\operatorname{tr}(C_{1}C_{2})=\operatorname{tr}(C_{1}^{1/2}C_{2}C_{1}^{1/2})\geq\mu\operatorname{tr}(C_{1})$
-holds. $\Box$
+Since $C_{2}\succeq\mu I_{\ell}$, we obtain $C_{1}^{1/2}C_{2}C_{1}^{1/2}\succeq\mu C_{1}$.
+Therefore, $\operatorname{tr}(C_{1}C_{2})=\operatorname{tr}(C_{1}^{1/2}C_{2}C_{1}^{1/2})\geq\mu\operatorname{tr}(C_{1})$.
+$\Box$
 Then, we can prove Theorems 2 and 3.
@@
 Similarly, we can prove that $h^{(k)}$ is lower semicontinuous.
-The proof proceeds in the same manner as that of Theorem 2.
+As $T\to\infty$, we have $D(T)^{-1}\to\begin{pmatrix}I_{n_{-}}\\ &O\end{pmatrix}$, and Theorem 1 implies that $D(T_{k})^{-*}W(\varphi(\widetilde{p});T_{k})^{-1}D(T_{k})^{-1}\to\begin{pmatrix}W_{-}(\varphi(\widetilde{p}))^{-1}\\ &O\end{pmatrix}$.
+Therefore, the remainder of the proof follows the same steps as in the proof of Theorem 2.
 $\Box$
 Appendix D Proof of Theorems 4 and 5
-To prove Theorems 4 and 5, we prepare the following lemmas, which can be proved in the same manner as [23, Theorems 1 and 3], although Problems 2 and 6 have slightly different forms.
+To prove Theorems 4 and 5, we prepare the following lemmas, which can be proved in the same manner as [24, Theorems 1 and 3], although Problems 2 and 6 have slightly different forms.
 Lemma 7
@@
 $\widetilde{X}_{0,0}^{(\ell)}\coloneq\begin{pmatrix}\widetilde{x}_{\alpha+1,\alpha+1}&\cdots&\widetilde{x}_{\alpha+1,\beta}\\ \vdots&\ddots&\vdots\\ \widetilde{x}_{\beta,\alpha+1}&\cdots&\widetilde{x}_{\beta,\beta}\end{pmatrix}\in\mathbb{C}^{m_{0}^{(\ell)}\times m_{0}^{(\ell)}},$
-where we set $\alpha=m_{-}+m_{0}^{(1)}+\cdots+m_{0}^{(\ell-1)},\ \beta=m_{-}+m_{0}^{(1)}+\cdots+m_{0}^{(\ell)}$ to simplify the indexing.
+where we set $\alpha=m_{-}+m_{0}^{(1)}+\cdots+m_{0}^{(\ell-1)},\ \beta=\alpha+m_{0}^{(\ell)}$ to simplify the indexing.
 Then, under the assumption that $W_{\infty}(x)=O$, it follows that $\widetilde{X}_{-,-}=O,\widetilde{X}_{+,+}=O$, and $\widetilde{X}_{0,0}^{(\ell)}=O\ (\ell=1,\ldots,m_{0}^{\prime})$.
@@
 Furthermore, we obtain
-$\widetilde{X}_{-,-}=\sum_{i=1}^{n}x_{i}q_{-,i}q_{-,i}^{*},$
-$\widetilde{X}_{+,+}=\sum_{i=1}^{n}x_{i}q_{+,i}q_{+,i}^{*},$
+$\widetilde{X}_{-,-}=\sum_{i=1}^{n}x_{i}q_{-,i}q_{-,i}^{*},\quad\widetilde{X}_{+,+}=\sum_{i=1}^{n}x_{i}q_{+,i}q_{+,i}^{*},$
 $\widetilde{X}_{0,0}^{(\ell)}=\sum_{i=1}^{n}x_{i}q_{0,i}^{(\ell)}q_{0,i}^{(\ell)}{}^{*}\quad(\ell=1,\ldots,m_{0}^{\prime}).$
@@
 Then, from Lemma 8, it suffices to prove $x=0$.
-Let $\widetilde{X}\coloneq Q^{-1}XQ^{-*}$ and partition $\widetilde{X}$ in the same way as 30.
+Let $\widetilde{X}\coloneq Q^{-1}XQ^{-*}$ and partition $\widetilde{X}$ in the same way as 31.
 By an argument analogous to that of Theorem 4, we obtain $\widetilde{X}_{-,-}=O$.
@@
 For Hurwitz matrices, it has been shown that the optimal solution is unique for all $T>0$.
-Proposition 4 ([23, Theorem 4])
+Proposition 4 ([24, Theorem 4])
 Assume that $A$ is Hurwitz. Then, Problems 1 and 2 each admit a unique optimal solution for all $T>0$.
@@
 For graph Laplacians of undirected connected graphs, the following result has been established.
-Proposition 5 ([23, Theorem 5])
+Proposition 5 ([24, Theorem 5])
 Let $L$ be the Laplacian matrix of an undirected connected graph, and let A=LA=-L. Then, Problems 1 and 2 each admit a unique optimal solution for all $T>0$.
 Without restricting the class of systems, the following result has been established for arbitrary matrices.
-Proposition 6 ([22, Theorem 1])
+Proposition 6 ([23, Theorem 1])
 Let $A$ be an arbitrary matrix. Then, Problems 1 and 2 each admit a unique optimal solution for almost all $T>0$.
@@
 Theorem 7
-Let
+Let $\lambda_{i}\ (i=1,\ldots,m)$ be the eigenvalues of $A$ as in Section III-A,
 $\Theta\coloneq\left\{\theta\in\mathbb{R}\setminus\{0\}\,\middle|\,\text{$\lambda_{i}+\lambda_{j}=\theta\sqrt{-1}$ for some $i,j$}\right\},$
@@
 Let $f\colon\mathbb{C}\to\mathbb{C}$ be a holomorphic function. Then, every eigenvalue of $f(K)$ is of the form $f(\mu)$, where $\mu$ is an eigenvalue of $K$.
-Moreover, we employ the following lemma, which can be proved in the same way as in [23, Theorems 1 and 3].
+Moreover, we employ the following lemma, which can be proved in the same way as in [24, Theorems 1 and 3].
 Lemma 11
@@
 We therefore investigate whether $F_{T}$ has zero as an eigenvalue.
-It follows from 31 that
-$\operatorname{vec}\left(\widetilde{W}(B;T)\right)$ $=\int_{0}^{T}\operatorname{vec}\left(\mathrm{e}^{At}B\mathrm{e}^{A^{\top}t}\right)\mathrm{d}t$
-$=\left(\int_{0}^{T}\mathrm{e}^{At}\otimes\mathrm{e}^{At}\mathrm{d}t\right)\operatorname{vec}(B)$
-$=\left(\int_{0}^{T}\mathrm{e}^{Kt}\mathrm{d}t\right)\operatorname{vec}(B).$
+It follows from 32 that
+$\operatorname{vec}\left(\widetilde{W}(B;T)\right)=\int_{0}^{T}\operatorname{vec}\left(\mathrm{e}^{At}B\mathrm{e}^{A^{\top}t}\right)\mathrm{d}t\\ =\left(\int_{0}^{T}\mathrm{e}^{At}\otimes\mathrm{e}^{At}\mathrm{d}t\right)\operatorname{vec}(B)=\left(\int_{0}^{T}\mathrm{e}^{Kt}\mathrm{d}t\right)\operatorname{vec}(B).$
 Thus, by vectorization, the map $F_{T}$ can be identified with the linear map
@@
 which is a holomorphic function.
-By identifying 32 and 33, we have $F_{T}=f_{T}(K)$.
+By identifying 33 and 34, we have $F_{T}=f_{T}(K)$.
 Consequently, from Lemmas 9 and 10, every eigenvalue of $F_{T}$ is of the form $f_{T}(\lambda_{i}+\lambda_{j})$ for some $i,j$.